### v2

# Sparse Multivariate Linear Regression with Strongly Associated Response Variables

Abstract
We propose new methods for multivariate linear regression when the regression coefficient matrix is sparse and the error covariance matrix is dense. We assume that the error covariance matrix has equicorrelation across the response variables. Two procedures are proposed: one is based on constant marginal response variance (compound symmetry), and the other is based on general varying marginal response variance. Two approximate procedures are also developed for high dimensions. We propose an approximation to the Gaussian validation likelihood for tuning parameter selection. Extensive numerical experiments illustrate when our procedures outperform relevant competitors as well as their robustness to model misspecification.
keywords:
High-dimensional multivariate linear regression, equicorrelation
1 Introduction
Multivariate linear regression simultaneously models multiple response variables in terms of one or more predictors. It is well studied, and we direct readers to Reinsel et al. [10] for a detailed review.
We focus on fitting multivariate linear regression models by penalized or constrained Gaussian likelihood. One of the first approaches maximizes the Gaussian likelihood subject to a rank constraint on the regression coefficient matrix [5, 10]. Like other dimension reduction methods, interpreting the fitted model in terms of the original predictors may be difficult.
Rothman et al. [12] proposed to jointly estimate the error precision matrix and the regression coefficient matrix by penalized Gaussian likelihood. They used $L_{1}$-penalties to encourage sparsity in estimates of the regression coefficients and error precision matrix, which can lead to easy-to-interpret fitted models. They showed that using the Gaussian loglikelihood, which accounts for the association between the response components, leads to better parameter estimation than using a multivariate residual sum of squares criterion, which does not account for this association.
Similar approaches that jointly estimate the error precision matrix and the regression coefficient matrix have been proposed. For instance, Lee and Liu [6] assumed that the response variables and the predictors have a joint multivariate normal distribution, and Wang [14] decomposed the multivariate regression problem into a series of penalized conditional log-likelihood of each response conditional on the predictors and other responses. Navon and Rosset [7] extended this to multivariate linear mixed effects models, and Chang and Welsh [3] considered an extension of Rothman et al. [12]’s method to multivariate robust linear regression. Zhu [17] considered a convex reparametrization of Rothman et al. [12]’s joint optimization problem. These methods assumed that the error precision matrix is sparse, which may be unreasonable in some applications with highly correlated response components. In addition, in high-dimensional settings, the graphical lasso subproblem [4] used in the block-wise coordinate decent algorithm of Rothman et al. [12] struggles when the error precision matrix is dense. These dense cases call for values of the penalty tuning parameter near zero, which lead to algorithm failures or very long computing times when solving the graphical lasso subproblem.
This motivated us to develop a new multivariate linear regression method for high-dimensional settings that works when the error precision matrix is dense. Similarly to Rothman et al. [12], we jointly estimate the regression coefficient matrix and the error precision matrix by minimizing the negative Gaussian loglikelihood. However, unlike their approach, we assume the error covariance matrix has compound symmetry (or equicorrelation). Our primary goal is to estimate the regression coefficient matrix when the responses are highly correlated. Although our assumed error correlation structure is simple, our estimation of the regression coefficient matrix is robust to misspecification of the error correlation. We propose an efficient computational algorithm to compute our estimators and we also propose an approximation to the validation likelihood for tuning parameter selection.
2 Problem setup and proposed estimator
Let $x_{i}=(x_{i1},\ldots,x_{ip})^{T}$ be the $p$-dimensional vector of nonrandom predictor values; let $y_{i}=(y_{i1},\ldots,y_{iq})^{T}$ be the observed $q$-dimensional response; and let $\epsilon_{i}=(\epsilon_{i1},\ldots,\epsilon_{iq})^{T}$ be the error for the $i$th subject $(i=1,\ldots,n)$. The multivariate linear regression model assumes that $y_{i}$ is a realization of
$Y_{i}=B_{*}^{T}x_{i}+\epsilon_{i},\qquad\quad i=1,\ldots,n,$ (1)
where $B_{*}\in\mathbb{R}^{p\times q}$ is the unknown regression coefficient matrix; and $\epsilon_{1},\ldots,\epsilon_{n}$ are iid $N_{q}(0,\Sigma_{*})$. To allow for highly correlated response variables, we assume that
$\Sigma_{*}=\eta^{2}_{*}\left\{(1-\theta_{*})I_{q}+\theta_{*}1_{q}1_{q}^{T}\right\},$ (2)
where $\eta_{*}\in(0,\infty)$ and $\theta_{*}\in[0,1)$ are unknown. This implies that ${\rm var}(\epsilon_{i,j})=\eta^{2}_{*}$ for $(i,j)\in\{1,\ldots,n\}\times\{1,\ldots,q\}$ and that ${\rm corr}(\epsilon_{i,j},\epsilon_{i,k})=\theta_{*}$ when $j\neq k$. This implicitly assumes that all of the response components are on the same scale. When they are not, we propose a different method described in Section 4. Our numerical experiments suggest that our estimation of $B_{*}$, which is our primary target, is robust to misspecification of $\Sigma_{*}$.
To write the negative loglikelihood, we first express (1) in terms of matrices: let $Y\in\mathbb{R}^{n\times q}$ have $i$th row $Y_{i}^{T}$; let $X\in\mathbb{R}^{n\times p}$ have $i$th row $x_{i}^{T}$; and let $E\in\mathbb{R}^{n\times q}$ have $i$th row $\epsilon_{i}^{T}$. Then (1) is $Y=XB_{*}+E$. The negative log-likelihood function evaluated at $(B,\Omega)$, where $\Omega$ is the variable corresponding to the inverse of $\Sigma_{*}$, is
$\mathcal{L}(B,\Omega)={\rm tr}\left[\frac{1}{n}(Y-XB)^{T}(Y-XB)\Omega\right]-\log|\Omega|.$
We propose the following penalized likelihood estimator of $(B_{*},\eta_{*}^{2},\theta_{*})$:
$(\hat{B},\hat{\eta}^{2},\hat{\theta})=\mathop{arg\,min}\displaylimits_{(B,\eta^{2},\theta)\in\mathbb{R}^{p\times q}\times(0,\infty)\times[0,1)}\left\{\mathcal{L}\left(B,\Omega(\eta^{2},\theta)\right)+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|\right\},$ (3)
where $\Omega(\eta^{2},\theta)=\left[\eta^{2}\left\{\theta 1_{q}1_{q}^{\prime}+(1-\theta)I_{q}\right\}\right]^{-1}$ and $\lambda\in[0,\infty)$ is a tuning parameter. Similarly to Rothman et al. [12], the $L_{1}$ penalty on $B$ encourages sparse regression coefficient estimation. We label the solution to this optimization as the Multivariate Regression with Compound Symmetry [MRCS] estimator. When the response components are on different scales, we define an alternative estimator described in Section 4.
Our goal is to estimate $B_{*}$ and predict future response values: we expect that the assumed error covariance structure in (2) will be an oversimplification in many applications. When $\lambda=0$, it is known that the regression coefficient estimator that solves (3) would be unchanged if $\Omega(\eta^{2},\theta)$ were replaced by any positive definite matrix [12]. So the assumed error correlation structure only influences how the entries in the regression coefficient matrix estimator are shrunk towards zero when $\lambda>0$. Our numerical experiments suggest that our estimation of $B_{*}$ is robust to misspecification of the error covariance.
3 Computational algorithms
3.1 Algorithm for exact computation
By the derivation illustrated in A of our supplementary material, the penalized estimator in (3) is equivalent to the following:
$\begin{split}\mathop{arg\,min}\displaylimits_{(B,\eta^{2},\theta)\in\mathbb{R}^{p\times q}\times(0,\infty)\times[0,1)}F_{\lambda}(B,\eta^{2},\theta;Y,X),\end{split}$ (4)
where
$\begin{split}F_{\lambda}(B,\eta^{2},\theta;Y,X)&=\frac{1}{n\eta^{2}(1-\theta)}\|Y-XB\|_{F}^{2}-\frac{\theta}{n\eta^{2}(1-\theta)(1-\theta+q\theta)}\|(Y-XB)1_{q}\|^{2}\\ &\indent+(q-1)\log(1-\theta)+\log(1+\{q-1\}\theta)+q\log(\eta^{2})\\ &\indent+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|.\end{split}$
We solve (4) by blockwise coordinate descent. We treat $(\eta^{2},\theta)$ as the first block, and $B$ as the second block. We let the superscript $(k)$ denote the $k$th iterate of each component that is being updated.
For a fixed $\hat{B}^{(k)}$, we reparametrize the problem via $\alpha=\eta^{2}(1-\theta),\,\gamma=\eta^{2}(1+(q-1)\theta)$, which is a one-to-one mapping from $(\eta^{2},\theta)$. Then (4) with $\hat{B}^{(k)}$ fixed becomes
$(\hat{\alpha}^{(k+1)},\hat{\gamma}^{(k+1)})=\mathop{arg\,min}\displaylimits_{0<\alpha\leq\gamma<\infty}\left\{\frac{M_{1}^{(k)}}{\alpha}-\frac{1}{q}\left(\frac{1}{\alpha}-\frac{1}{\gamma}\right)M_{2}^{(k)}+\log(\gamma)+(q-1)\log(\alpha)\right\},$ (5)
where $M_{1}^{(k)}=\frac{1}{n}\|Y-X\hat{B}^{(k)}\|_{F}^{2}$ and $M_{2}^{(k)}=\frac{1}{n}\|(Y-X\hat{B}^{(k)})1_{q}\|^{2}$. Since the objective function in (5) is separable, by first order condition, we get
$\hat{\alpha}^{(k+1)}=\frac{qM_{1}^{(k)}-M_{2}^{(k)}}{q(q-1)},$ (6)
$\hat{\gamma}^{(k+1)}=\hat{\gamma}^{(k+1)}(\hat{\alpha}^{(k+1)})=\max\left\{\hat{\alpha}^{(k+1)},\frac{M_{2}^{(k)}}{q}\right\},$ (7)
where $M_{1}^{(k)}\geq M_{2}^{(k)}/q$ holds by the Cauchy-Schwarz inequality. Finally,
$((\hat{\eta}^{2})^{(k)},\hat{\theta}^{(k)})=\left(\hat{\alpha}^{(k)}+\frac{\hat{\gamma}^{(k)}-\hat{\alpha}^{(k)}}{q},\frac{\hat{\gamma}^{(k)}-\hat{\alpha}^{(k)}}{\hat{\gamma}^{(k)}+(q-1)\hat{\alpha}^{(k)}}\right).$ (8)
Now, when $\hat{\Omega}^{(k+1)}((\hat{\eta}^{2})^{(k+1)},\hat{\theta}^{(k+1)})$ is fixed, the corresponding optimization problem is given by
$\hat{B}^{(k+1)}(\hat{\Omega}^{(k+1)})=\mathop{arg\,min}\displaylimits_{B}\left\{{\rm tr}\left[\frac{1}{n}(Y-XB)^{T}(Y-XB)\hat{\Omega}^{(k+1)}\right]\right\}+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|.$ (9)
We solve (9) through a function rblasso in the R package MRCE [11].
To improve the performance with a better initializer, we suggest the following procedure. We first perform $q$ lasso regressions for each response with the same optimal tuning parameter $\hat{\lambda}_{0}$ selected with a cross validation. Let $\hat{B}_{\hat{\lambda}_{0}}$ be the solution. We refer this as combined lasso initializer. We initialize the algorithm from $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. We set the convergence tolerance parameter $\varepsilon=10^{-7}$. We summarize the algorithmic procedure in Algorithm 1 below. Steps 1 and 2 both guarantee a decrease in the objective function value.
Algorithm 1 Multivariate Regression with Compound Symmetry [MRCS]
  For each fixed value of $\lambda$, initialize $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Set $k=0$ and $((\hat{\eta}^{2})^{(0)},\hat{\theta}^{(0)})=(1,0)$.
  Step 1: Compute $((\hat{\eta}^{2})^{(k+1)},\hat{\theta}^{(k+1)})(\hat{B}^{(k)})$ by (8).
  Step 2: Compute $\hat{B}^{(k+1)}((\hat{\eta}^{2})^{(k+1)},\hat{\theta}^{(k+1)})$ by (9).
  Step 3: If $|F_{\lambda}(\hat{B}^{(k+1)},(\hat{\eta}^{2})^{(k+1)},\hat{\theta}^{(k+1)})-F_{\lambda}(\hat{B}^{(k)},(\hat{\eta}^{2})^{(k)},\hat{\theta}^{(k)})|<\varepsilon\operatorname{tr}(Y^{\prime}Y)/n$ then stop. Otherwise go to Step 1 and $k\leftarrow k+1$.
3.2 Algorithm for approximate computation
When $p\geq n$, the residual sample covariance matrix, $(Y-XB)^{T}(Y-X)/n$, is singular. When there exists $\bar{B}$ such that any one of the columns of $Y-X\bar{B}$ is zero, then the objective function value for updating $\Omega$ can be made arbitrarily negative by increasing the corresponding diagonal element of $\Omega$ sufficiently, which implies that a global minimizer does not exist. In addition, the computational time of the blockwise coordinate descent for fitting both regression coefficients and the error precision matrix tends to be significantly higher in increased dimensions of $p,\,q$ [12]. To improve computational efficiency and ensure the well-posed optimization problem in high dimensions, we provide an approximate solution to our procedure by following the idea of approximate MRCE proposed in Rothman et al. [12]. As in Algorithm 1, we compute the initial $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Then, for each $\lambda$, we compute $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ by (8). After this step for (inverse) covariance update, we compute the proposed approximate solution $\hat{B}$ with known $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ by (9). We suppress the dependence of $\hat{B}$ and $(\hat{\eta}^{2}(\hat{B}_{\hat{\lambda}_{0}}),\hat{\theta}(\hat{B}_{\hat{\lambda}_{0}}))$ on $\lambda$ for notational simplicity. We refer the approximate solution to (4) as ap.MRCS. We summarize the procedure in Algorithm 2 below.
Algorithm 2 Approximated version of Multivariate Regression with Compound Symmetry [ap.MRCS]
  For each fixed value of $\lambda$, initialize $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$.
  Step 1: Compute $(\hat{\eta}^{2},\hat{\theta})(\hat{B}^{(0)})$ by (8).
  Step 2: Compute $\hat{B}(\hat{\eta}^{2},\hat{\theta})$ by (9).
4 Extension to general equicorrelation covariance matrix
When the response components are on different scales, then it may be unreasonable to assume that the components of the errors have the same variance $\eta_{*}^{2}$. We extend our model by allowing different $\eta_{*j}$’s for each response component. The model follows (1) except that we use the following error covariance matrix:
$\Sigma_{*}$ $={\rm diag}(\{\eta_{*j}\}_{j=1}^{q})\left[(1-\theta)I_{q}+\theta 1_{q}1_{q}^{T}\right]{\rm diag}(\{\eta_{*j}\}_{j=1}^{q}),$
where $(\eta_{*1},\ldots,\eta_{*q})\in(0,\infty)^{q}$ are the unknown standard deviations of the $q$ marginal distributions of the error, i.e. ${\rm var}(\epsilon_{i,j})=\eta_{*j}^{2}$ for $(i,j)\in\{1,\ldots,n\}\times\{1,\ldots,q\}$. The inverse error covariance matrix is
$\Omega_{*}=\frac{1}{1-\theta_{*}}{\rm diag}(\{\eta_{*i}^{-1}\}_{i=1}^{q})\left[I_{q}-\frac{\theta_{*}}{1+(q-1)\theta_{*}}1_{q}1_{q}^{T}\right]{\rm diag}(\{\eta_{*i}^{-1}\}_{i=1}^{q}).$ (10)
Then the corresponding penalized negative log-likelihood optimization is
$\begin{split}\mathop{arg\,min}\displaylimits_{(B,\{\eta_{i}\}_{i=1}^{q},\theta)\in\mathbb{R}^{p\times q}\times\mathbb{R}_{+}^{q}\times[0,1)}F^{gen}_{\lambda}(B,\{\eta_{i}\}_{i=1}^{q},\theta;Y,X),\end{split}$ (11)
where
$\begin{split}F^{gen}_{\lambda}(B,\{\eta_{i}\}_{i=1}^{q},\theta;Y,X)&=\frac{1}{n(1-\theta)}\|(Y-XB){\rm diag}(\{\eta_{i}^{-1}\})\|_{F}^{2}\\ &\indent-\frac{\theta}{n(1-\theta)(1-\theta+q\theta)}\|[(Y-XB){\rm diag}(\{\eta_{i}^{-1}\})]1_{q}\|^{2}\\ &\indent+(q-1)\log(1-\theta)+\log(1+\{q-1\}\theta)+2\sum_{i=1}^{q}\log(\eta_{i})\\ &\indent+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|,\end{split}$
and $\lambda\in[0,\infty)$ is a tuning parameter. Note that the problem (11) is bi-convex in $B$ and ($(\{\eta_{i}\}_{i=1}^{q},\theta)$. As in Algorithm 3.2, we treat ($(\{\eta_{i}\}_{i=1}^{q},\theta)$ as the first block and $B$ as the second block in blockwise coordinate descent.
For a fixed $\hat{B}^{(k)}$, we solve for ($(\{\eta_{i}\}_{i=1}^{q},\theta)$ cyclically. By simple algebra (see A in our supplementary material), we compute the update for $\eta_{j}$ by
$\hat{\eta}_{j}^{(k+1)}=\frac{-K_{1}^{(k)}+\sqrt{(K_{1}^{(k)})^{2}+4K^{(k)}_{2}}}{2},$ (12)
where
$K_{1}^{(k)}=\frac{\hat{\theta}^{(k)}}{n(1-\hat{\theta}^{(k)})(1+(q-1)\hat{\theta}^{(k)})}\sum_{i=1}^{n}e_{ij}^{(k)}\left(\sum_{k\neq j}e_{ik}^{(k)}\frac{1}{\hat{\eta}^{(k)}_{k}}\right),$
$K_{2}^{(k)}=\frac{1+(q-2)\hat{\theta}^{(k)}}{n(1+(q-1)\hat{\theta}^{(k)})(1-\hat{\theta}^{(k)})}\left(\sum_{i=1}^{n}(e_{ij}^{(k)})^{2}\right),$
and $e_{ij}^{(k)}$ is the $(i,j)$-th element of $Y-X\hat{B}^{(k)}$. After sequentially solving for $\{\eta_{i}\}_{i=1}^{q}$, we solve the following by line search:
$\begin{split}\hat{\theta}^{(k+1)}=\mathop{arg\,min}\displaylimits_{\theta\in[0,1)}&\frac{1}{n(1-\theta)}\|(Y-X\hat{B}^{(k)}){\rm diag}(\{(\hat{\eta}_{i}^{(k+1)})^{-1}\})\|_{F}^{2}\\ &\indent-\frac{\theta}{n(1-\theta)(1-\theta+q\theta)}\|[(Y-X\hat{B}^{(k)}){\rm diag}(\{(\hat{\eta}_{i}^{(k+1)})^{-1}\})]1_{q}\|^{2}\\ &\indent+(q-1)\log(1-\theta)+\log(1+\{q-1\}\theta).\end{split}$ (13)
For efficiency, we do not require full convergence in each inner loop over ($(\{\eta_{i}\}_{i=1}^{q},\theta)$. Instead, we conduct single-iteration updates for each $\{\eta_{i}\}_{i=1}^{q}$ and $\theta$ by (12) and (13).
We compute $\hat{B}^{(k+1)}$ using (9) with $\hat{\Omega}^{(k+1)}$ set to the right side of (10) with $\{\eta_{i}\}_{i=1}^{q}$ and $\theta$ replaced by $\{\hat{\eta}_{i}^{(k+1)}\}_{i=1}^{q}$ and $\hat{\theta}^{(k+1)}$, respectively. We call the solution to (11) MRGCS.
We again start the algorithm with combined lasso initializer as in Algorithm 1. We compared the combined lasso initializer with a separate lasso initializer that fits $q$ separate lasso regressions for each response with the different tuning parameters selected by cross validation for each response. However, we found that the combined lasso initializer generally performed better than the separate lasso except for a few cases. We again set the convergence tolerance parameter $\varepsilon=10^{-7}$ for the following Algorithm 3 which summarizes the entire procedure.
Algorithm 3 Multivariate Regression with Generalized Compound Symmetry [MRGCS]
  For each fixed value of $\lambda$, initialize $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$. Set $k=0$ and ($(\{\hat{\eta}_{i}^{(0)}\}_{i=1}^{q},\hat{\theta}^{(0)})=(1_{q},0)$.
  Step 1: Compute one-step update of ($(\{\hat{\eta}_{i}^{(k+1)}\}_{i=1}^{q},\hat{\theta}^{(k+1)})(\hat{B}^{(k)})$ by solving (12)–(13) cyclically.
  Step 2: Compute $\hat{B}^{(k+1)}(\{\hat{\eta}_{i}^{(k+1)}\}_{i=1}^{q},\hat{\theta}^{(k+1)})$ by solving (9).
  Step 3: If $|F^{gen}_{\lambda}(\hat{B}^{(k+1)},\{\hat{\eta}_{i}^{(k+1)}\}_{i=1}^{q},\hat{\theta}^{(k+1)})-F^{gen}_{\lambda}(\hat{B}^{(k)},\{\hat{\eta}_{i}^{(k)}\}_{i=1}^{q},\hat{\theta}^{(k)})|<\varepsilon\operatorname{tr}(Y^{\prime}Y)/n$ then stop. Otherwise go to Step 1 with $k\leftarrow k+1$.
4.1 Approximate solutions II
As in Section 3.2, we propose an approximate solution to (11). Again, we start the algorithm with $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$, the combined lasso initializer. Then, for each $\lambda$, we compute ($(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})$ from (11). However, the difference between this approximation and the canonical MRGCS is that we conduct the updating step for ($(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})$ until convergence instead of a single-iteration update. This subproblem is obviously convex in ($(\{\eta_{i}\}_{i=1}^{q},\theta)$. Next, we compute the final $\hat{B}$ by solving (9) once. We call this approximation ap.MRGCS. We summarize it in Algorithm 4 below.
Algorithm 4 Approximated version of Multivariate Regression with generalized Compound Symmetry [ap.MRGCS]
  For each fixed value of $\lambda$, initialize $\hat{B}^{(0)}=\hat{B}_{\hat{\lambda}_{0}}$.
  Step 1: Compute ($(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})(\hat{B}^{(0)})$ cyclically by solving (12) and (13) until convergence.
  Step 2: Compute $\hat{B}(\{\hat{\eta}_{i}\}_{i=1}^{q},\hat{\theta})$ by solving (9).
5 Asymptotic statistical properties
We study large sample asymptotic behavior of our estimator MRCS defined by (4) using the same asymptotic framework as Lee and Liu [6]. We study high-dimensional statistical properties of ap.MRCS (see Algorithm 3.2) in C of the supplementary material. We keep $p$ and $q$ fixed with $n>p+q$ throughout this section. As stated in Proposition 1 of Zou [18], the equally-weighted lasso-penalized estimator does not possess the oracle property. Thus, to ensure the oracle property, it is encouraged to assign weights, $w_{jk}$, to each $|B_{jk}|$ in the penalty term. We let $w_{jk}=1/|B^{({\rm ols})}_{jk}|^{r}$ ($r>1$) be the weight for $|B_{jk}|$, where $B^{({\rm ols})}_{jk}$ is $(j,k)$-th element of the regression coefficient matrix obtained from ordinary least squares. Using these weights, the slightly modified optimization (than (4)) we consider in this section is as follows:
$\begin{split}\mathop{arg\,min}\displaylimits_{(B,\eta^{2},\theta)\in\mathbb{R}^{p\times q}\times(0,\infty)\times[0,1)}&\frac{1}{\eta^{2}(1-\theta)}\operatorname{tr}\left\{n^{-1}(Y-XB)^{T}(Y-XB)\left[I_{q}-\frac{\theta}{1-\theta+q\theta}1_{q}1^{T}_{q}\right]\right\}\\ &\quad+(q-1)\log(1-\theta)+\log(1+\{q-1\}\theta)+q\log(\eta^{2})\\ &\quad+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}w_{jk}|B_{jk}|.\end{split}$ (14)
We make the following assumptions to state large sample asymptotics of MRCS.
Assumption A.
blah
  1. (A1)
    : $X^{T}X/n\rightarrow Z$, where $Z$ is a positive definite matrix.
  2. (A2)
    : There exists $\tilde{B}_{jk}$ a $\sqrt{n}$-consistent estimator of $B_{jk*}$, where $B_{jk*}$ is the $(j,k)$-th element of $B_{*}$ for $(j,k)\in\{1,\ldots,p\}\times\{1,\ldots,q\}$.
  3. (A3)
    : There exists $(\tilde{\eta}^{2},\tilde{\theta})$ $\sqrt{n}$-consistent estimator of $(\eta_{*}^{2},\theta_{*})$.
  4. (A4)
    : The distribution of $E$ has finite joint fourth moments.
Condition (A1) is a standard condition in linear regression large sample asymptotics literature, and was also assumed in Zou [18], Lee and Liu [6], and Chang and Welsh [3]. It implies that the design matrix has good asymptotic behavior. We note that (A2) and (A3) are generally satisfied by maximum likelihood estimators [MLEs] or $L_{2}$-penalized MLEs. (A4) is satisfied by broad class of error distributions such as multivariate sub-exponential distributions.
For a matrix $R\in\mathbb{R}^{p\times q}$, we let ${\rm vec}(R)\in\mathbb{R}^{pq}$ be the vector formed by stacking the columns of $R$. Define $S:=\{i:{\rm vec}(B_{*})_{i}\neq 0\}$. Let $v_{A}$ be the subvector of the of the entries in $v$ with indices in $A$. For a square matrix $M\in\mathbb{R}^{q\times q}$, we further define $M_{A}$ as the $|A|\times|A|$ matrix obtained by removing the $i$-th row and column of M for $i\in A^{c}$, where $A\subseteq\{1,\ldots,q\}$. We let $\otimes$ denote the Kronecker product of two matrices. Recall that the inverse of compound symmetry error covariance is computed as
$\Omega_{*}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}\left[I_{q}-\frac{\theta_{*}}{1+(q-1)\theta_{*}}1_{q}1_{q}^{T}\right].$
In the following theorem, we provide the oracle guarantee, the limit distribution of $\hat{B}$, and the joint limit distribution of $(\hat{\eta}^{2},\hat{\theta})$. This result is analogous to Theorem 3 of Lee and Liu [6].
Theorem 1.
Under the conditions (A1)(A4), assuming that $\lambda\sqrt{n}\rightarrow 0$ and $\lambda n^{(r+1)/2}\rightarrow\infty$, then, there exists a local minimizer $(\hat{B},\hat{\eta}^{2},\hat{\theta})$ to the optimization problem (14) that satisfies
$\lim_{n\rightarrow\infty}P(\hat{B}_{jk}=0)=1\qquad\qquad{\rm if~{}}B_{jk*}=0,$
$\sqrt{n}({\rm vec}(\hat{B})_{S}-{\rm vec}(B_{*})_{S})\rightarrow_{d}N(0,D^{-1}),$
$\left\|(\hat{\eta}^{2},\hat{\theta})-(\eta_{*}^{2},\theta_{*})\right\|=O_{p}(1/\sqrt{n}),$
where $D=(\Omega_{*}\otimes Z)_{S}$. Furthermore, if $\theta_{*}\in(0,1)$, then
$\sqrt{n}((\hat{\eta}^{2},\hat{\theta})-(\eta_{*}^{2},\theta_{*}))\rightarrow_{d}W^{T}N_{2}(0,V),$
where $W=(1,-1/\eta_{*}^{2})^{T}$, and $V$ is defined in B of the supplementary material (see (27)).
6 Tuning parameter selection
We suggest the following tuning parameter selection procedure for MRCS, ap.MRCS, MRGCS, and ap.MRGCS:
$\hat{\lambda}=\mathop{arg\,min}\displaylimits_{\lambda\in\Lambda}\sum_{k=1}^{K}\operatorname{tr}\left\{n^{-1}(Y_{k}-X_{k}\hat{B}_{-k})^{T}(Y_{k}-X_{k}\hat{B}_{-k})\hat{\Omega}_{-k}\right\},$ (15)
where
$\hat{\Omega}_{-k}=\left[\hat{\eta}_{-k}\left\{\hat{\theta}_{-k}1_{q}1_{q}^{T}+(1-\hat{\theta}_{-k})I_{q}\right\}\right]^{-1}.$
In (15), $\Lambda$ is a candidate set of tuning parameter values, $K$ is the number of folds used for the cross validation, $\hat{B}_{-k}$ is the estimated regression coefficient matrix from each method based on the observations that excludes the $k$th fold, $Y_{k},\,X_{k}$ are the responses and predictors corresponding to $k$th fold, and $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ are the estimated parameters based on the observations that excludes the $k$th fold. The loss function in (15) is the negative Gaussian validation loglikelihood without the log determinant term. Including this term was slightly less stable than excluding it in our numerical experiments. Similar Gaussian validation likelihood minimization is presented in Lee and Liu [6], which used this loss criterion for finding an optimal tuning parameter associated graphical lasso applied to the error precision.
We note that we do not use $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ from the output of MRCS. Rather we use the one-step estimators used for ap.MRCS that are computed in the Step 1 of Algorithm 3.2. We use the same $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ for both MRCS and ap.MRCS. This is because the stability of ap.MRCS output turned out to be significantly better than that of the canonical MRCS. Furthermore, we use $\hat{\eta}_{-k},\,\hat{\theta}_{-k}$ which assumes compound symmetry error covariance even when we fit MRGCS or ap.MRGCS which assume varying marginal error variances. This is because the output of ap.MRCS was significantly more stable in the tuning parameter selection compared with that of ap.MRGCS as well as the MRGCS output. We provide further support regarding this tuning parameter selection procedure through extensive simulations in G.1 and G.2 of the supplementary material.
7 Simulations
The data for our simulations are generated from the following model:
$Y_{i}=\mu+X_{i}B_{*}+\epsilon_{i},\quad X_{i}\sim N(0,\Sigma_{X})\in\mathbb{R}^{p},\quad\epsilon_{i}\sim N(0,\Sigma_{*})\in\mathbb{R}^{q},\quad i=1,\ldots,n$ (16)
where $\mu=(1,1+4/(q-1),\ldots,5)$, $(\Sigma_{X})_{i,j}=(0.7)^{|i-j|}$ for $1\leq i,j\leq p$, and
$\Sigma_{*}={\rm diag}(\{\eta_{k}\}_{k=1}^{q})\left[(1-\theta)I_{q}+\theta 1_{q}1_{q}^{T}\right]{\rm diag}(\{\eta_{k}\}_{k=1}^{q}).$ (17)
We use $n=50$ training observations throughout. To generate $B_{*}$, we follow the same regression coefficient matrix generating procedure used in Section 3 of Rothman et al. [12]. We define the operation * as the element-wise matrix product. The coefficients matrix $B_{*}$ is generated by
$B_{*}=W*K*Q,$ (18)
where $W$ has iid entries from $N(0,1)$; $K$ has entries from iid $Ber(s_{1})$ (a Bernoulli distribution which returns $1$ with probability $s_{1}$); $Q=1_{p}1_{q}^{T}Q_{1}$ where a $q\times q$ diagonal matrix $Q_{1}$ has diagonal elements from iid $Ber(s_{2})$. This results in, $Q$ has rows that are either all one or all zero, where the population proportion of having all-one row ($1_{q}^{\prime}$) equals to $s_{2}$. Under this setting each model is expected to have $(1-s_{2})p$ predictors that are irrelevant for all $q$ responses, and each relevant predictor is expected to have contribution to $s_{1}q$ of the response variables.
We compare a separate lasso regression with a uniquely selected tuning parameter for each response [LASP], combined lasso [LAS] which employs the same tuning parameter for all responses, MRCE, approximate MRCE [ap.MRCE] [12], MRCS (the solution to (4)), MRGCS (the solution to (11)). In addition to the canonical estimators, MRCS and MRGCS, we also consider their approximate versions: ap.MRCS (see Section 3.2), as well as ap.MRGCS (see Section 4.1). Moreover, we compare these methods to an oracle procedure that assumes the true error precision $\Omega_{*}=\Sigma_{*}^{-1}$ is known, and only estimates $B$ by $L_{1}$-penalty.
$\begin{split}\hat{B}_{\rm Or}:=\mathop{arg\,min}\displaylimits_{B\in\mathbb{R}^{p\times q}}&\operatorname{tr}\left\{n^{-1}(Y-XB)^{T}(Y-XB)\Omega_{*}\right\}+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|.\end{split}$ (19)
We refer this estimator as MRCS-Or. In MRCE, we penalized the diagonals of the inverse covariance matrix only when $p\geq n$.
Tuning parameters are selected using 5 fold cross validation from $\lambda\in\Lambda$, where $\Lambda=\{10^{-4+0.5k}:k=0,1,\ldots,14\}$. The optimal tuning parameter selection procedures for our methods are discussed in Section 6 (see (15)). The optimal tuning parameter for MRCS-Or is also selected by (15) except that we use $\Omega_{*}$ in place of $\hat{\Omega}_{-k}$ in (15). For the other competitors, we select each tuning parameter that minimizes validation prediction error with 5 fold cross validation. For the estimators which require a single tuning parameter, we again use the candidate set as $\lambda\in\{10^{-4+0.5k}:k=0,1,\ldots,14\}$. And for MRCE, we use $(\lambda_{1},\lambda_{2})\in\Lambda_{1}\times\Lambda_{2}$, where $\Lambda_{1}=\Lambda_{2}=\{10^{-4+0.5k}:k=0,1,\ldots,14\}$.
For the primary criterion for the model comparison, we measure the model error, ${\rm tr}[(\hat{B}-B_{*})^{T}\Sigma_{X}(\hat{B}-B_{*})]$ [1, 16] with $\hat{B}$ provided by each method. We also measure the prediction error, $\|\hat{Y}-Y\|_{F}^{2}$, on the test set which has 200 observations which are generated from the same data generating process as in the training set. We further measure the true negative rate [TNR] and true positive rate [TPR] for the regression coefficient matrix estimation as follows [12]:
${\rm TNR}(\hat{B})=\frac{\#\{(i,j)\in[p]\times[q]:\hat{B}_{ij}=0,\,B_{ij*}=0\}}{\#\{(i,j)\in[p]\times[q]:B_{ij*}=0\}},$ (20)
${\rm TPR}(\hat{B})=\frac{\#\{(i,j)\in[p]\times[q]:\hat{B}_{ij}\neq 0,\,B_{ij*}\neq 0\}}{\#\{(i,j)\in[p]\times[q]:B_{ij*}\neq 0\}},$ (21)
where $[k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$; #$\#A$ stands for the number of elements in the set $A$.
7.1 Setting I: Constant $\eta$
In this section, we consider $(p,q)\in\{(20,50),(50,20),(80,80)\}$, We refer the readers to H.1 in the supplementary material for the results of low dimensional simulations ($p=q=20$). We vary $s_{1}\in\{0.1,0.5\},\,s_{2}\in\{0.1,0.5,1\},\theta\in\{0,0.5,0.75,0.9,0.95\}$, and fix $\eta_{i}=1$ for all $i=1,\ldots,q$. We drop MRCE due to its substantially high computation time. When $p\geq n$ we exclude MRCS, MRGCS from the set of competitors, where we still consider ap.MRCS, and ap.MRGCS, since they can avoid the residual covariance instability (see Section 3.2). The results for each setting are based on 50 independent replications. Throughout the results, the trend across the estimators in the prediction error is nearly equivalent to that in the model error. Hence, we mainly discuss the model error as our primary criterion.
7.1.1 Results when $(p,q)=(20,50)$
The complete simulation results are in Figure 1922. As $\theta$ increases, all our equicorrelation based estimators consistently outperforms ap.MRCE as well as two lasso methods. The best non-oracle estimators were MRCS and ap.MRCS. MRGCS and ap.MRGCS also showed competitive performance. In TNR, sep.lasso showed the best rate in general, MRCS and ap.MRCS outperformed MRGCS and ap.MRGCS, and ap.MRCE was the worst. On the other hand, in TPR, two lasso solutions were generally the worst, and all joint optimization methods showed increasing TPR with $\theta$. This indicates the joint optimization methods tend to provide more dense solution than lasso as $\theta$ gets increased.
7.1.2 Results when $(p,q)=(50,20)$
[…]
7.1.3 Results when $(p,q)=(80,80)$
The computing time for ap.MRCE is substantially higher than other procedures, specifically when the tuning parameter associated with the graphical lasso subproblem is close to $10^{-4}$. Given this issue we restricted the tuning parameter set for the $L_{1}$ penalty on the error precision into $\{10^{-2+0.5k}:k=0,1,\ldots,8\}$. We note that this computational setup for ap.MRCE is used for the other simulations where $(p,q)=(80,80)$ in Section 7.2 and H.2 (of the supplementary material).
The complete simulation results are in Figure 2730. The best non-oracle estimator was ap.MRCS. Again ap.MRGCS suffered only when $(s_{1},s_{2})=(0.5,0.1)$, and it showed competitive performance in the other settings. The general performance of ap.MRCE was significantly worse than the previous $(p,q)$ pairs.
7.2 Setting II: Equicorrelation of $\Sigma_{*}$ with heterogeneous and asymmetrically distributed $\eta_{i}$’s
In this section, we study the performance of our equicorrelation based estimators under generalized $\Sigma_{*}$ in which $\eta_{i}$’s are now heterogeneous. In H.2 of the supplementary material, we compared our methods to the others in the settings where $\eta_{i}$’s are still heterogeneous but symmetrically distributed. We discovered that MRCS and ap.MRCS showed comparable performance to MRGCS and ap.MRGCS, the two best non-oracle methods in that setting. This may be due to the symmetrically distributed $\eta_{i}$’s ($i\in\{1,\ldots,q\}$) which can smooth out heterogeneous marginal error variances. Now we consider asymmetric cases in this section. We used the same data generating process (16)–(18) except that we considered
$\begin{split}&\eta_{1},\ldots,\eta_{10}=1/2,\,\eta_{11},\ldots,\eta_{20}=1/\sqrt{2},\,\eta_{21},\ldots,\eta_{30}=1,\\ &\eta_{31},\ldots,\eta_{40}=\sqrt{3},\,\eta_{41},\ldots,\eta_{50}=3,\end{split}$ (22)
when $(p,q)=(20,50)$, and
$\begin{split}&\eta_{1},\ldots,\eta_{4}=1/2,\,\eta_{5},\ldots,\eta_{8}=1/\sqrt{2},\,\eta_{9},\ldots,\eta_{12}=1,\\ &\eta_{13},\ldots,\eta_{16}=\sqrt{3},\,\eta_{17},\ldots,\eta_{20}=3,\end{split}$ (23)
when $(p,q)=(50,20)$. Lastly, when $(p,q)=(80,80)$, we used
$\begin{split}&\eta_{1},\ldots,\eta_{10}=1/2,\,\eta_{11},\ldots,\eta_{20}=1/\sqrt{2},\,\eta_{21},\ldots,\eta_{30}=1/\sqrt[4]{2},\,\eta_{31},\ldots,\eta_{40}=1\\ &\eta_{41},\ldots,\eta_{50}=\sqrt{3},\,\eta_{51},\ldots,\eta_{65}=2,\,\eta_{66},\ldots,\eta_{80}=3.\end{split}$ (24)
To see the effect of varying $\eta_{i}$ on the regression coefficient shrinkage, we also compared the original $L_{1}$-penalty on $B$, $\lambda\sum_{j,k}|B_{jk}|$, with an adaptive $L_{1}$-penalty, $\lambda\sum_{j,k}|\eta_{k}^{-1}B_{jk}|$, for MRCS-Or. However, we did not find any supporting evidence on the use of the adaptive penalty instead of the original penalty. Thus, we still suggest the use of the original $L_{1}$-penalty function even in the case of varying $\eta_{i}$.
7.2.1 Results when $(p,q)=(20,50)$
The complete results can be found in Figure 3134. MRGCS and ap.MRGCS were the two best non-oracle methods, and consistently outperformed the other methods. Note that MRGCS was slightly better than its approximate counterpart. As opposed to the results shown in H.2.1 of the supplementary material where $\eta_{i}$’s are symmetrically distributed, MRCS and ap.MRCS struggled under the asymmetric setting discussed in this section. With relatively higher $s_{1}s_{2}$ values, ap.MRCE even outperformed these two methods for higher values of $\theta$. TNR/TPR results and trends are similar to those shown in Section 7.1.1.
7.2.2 Results when $(p,q)=(50,20)$
The complete results are in Figure 3538. ap.MRGCS was the best non-oracle method, although it suffered again when $(s_{1},s_{2})=(0.5,0.1)$. As opposed to the results in $(p,q)=(20,50)$, ap.MRCS was the second-best non-oracle competitor, and it outperformed ap.MRGCS when $s_{2}=0.1$. ap.MRCE showed poor performance generally. Note, the prediction performance of MRCS-Or, ap.MRGCS, and ap.MRCS was substantially better than the others. In TNR, surprisingly, ap.MRGCS was the worst, and ap.MRCS was significantly better than ap.MRGCS. In TPR, ap.MRCE and two lasso methods showed poor performance.
7.2.3 Results when $(p,q)=(80,80)$
The complete simulation results are in Figure 3942. The results show the best non-oracle estimator in this setting is ap.MRGCS. Again, ap.MRGCS suffered only when $(s_{1},s_{2})=(0.5,0.1)$ where it was outperformed by ap.MRCE as well as sep.lasso (see similar results in H.2.3 of the supplementary material where $\eta_{i}$’s are symmetrically distributed). ap.MRCS was outperformed by ap.MRCE for some values of $\theta$’s, when $(s_{1},s_{2})=(0.1,0.5),\,(0.1,1)$. Note that this can be compared with the results in H.2.3 of the supplementary material where ap.MRCS and ap.MRGCS substantially outperformed ap.MRCE. The relative performance of ap.MRCE was improved. ap.MRCE showed substantially lower TNR compared with the other procedures as $\theta$ gets increased. But, it was the best non-oracle method in TPR.
8 Simulation under the model misspecification
8.1 Misspecification of $B_{*}$: When the true $B_{*}$ is non-sparse
In this section, we study how non-sparse true regression coefficient matrix affects the proposed methods. As in previous simulations we measure the prediction and model error for evaluation. We follow the same data generating process (16)–(17). We consider $\eta_{i}$’s in (17) over three different settings: constant $\eta_{i}$’s (Setting A; Section 7.1), heterogeneous and symmetrically distributed $\eta_{i}$’s (Setting B; H.2), heterogeneous and asymmetrically distributed $\eta_{i}$’s (Setting C; Section 7.2). We do not consider different sparsity levels for $B_{*}$. Rather than using (18) for generating regression coefficients matrix $B_{*}$, for each element of $B_{*}$, we generate iid draws from a uniform distribution on $(-1/4,1/4)$ when $(p,q)=(20,50),\,(50,20)$, and on $(-1/10,1/10)$ when $(p,q)=(80,80)$. This regression coefficient matrix is non-sparse, and includes many non-zero, small signals. One can expect that sep.lasso and comb.lasso are likely to suffer from this design. We again set $\theta\in\{0,0.5,0.75,0.9,0.95\}$. In addition to the same set of competitors in Section 7, we also considered combined ridge [c.ridge], and separate ridge [s.ridge]. c.ridge and s.ridge are analogously defined as combined lasso and separate lasso.
The complete results are in Figure 4344. Under the setting A, we denote the sub-settings $(p,q)=(20,50)$, $(50,20)$, $(80,80)$ by A1, A2, and A3, respectively. And, we similarly denote the sub-settings for B and C. MRGCS and ap.MRGCS were the best non-oracle method in all settings, and outperformed ap.MRCE as well as lasso and ridge estimators as $\theta$ increases. We note that ap.MRCE was the third-best non-oracle method. In Setting A, the setting with constant $\eta_{i}$’s, MRCS and ap.MRCS were also regarded as the best non-oracles. Although they showed competitive performance with symmetric and heterogeneous $\eta_{i}$’s when $B_{*}$ has sparsity (see H.2 of the supplementary material), both methods struggled in non-sparse cases (settings B and C). The results from this section indicate that if non-sparse $B_{*}$ has sufficiently small (in absolute value) elements, then there is more gain by joint optimization than lose by misspecified $L_{1}$ penalty.
8.2 Misspecification of $\Sigma_{*}$: Corrupted compound symmetry to general non-sparse covariance
In this section, we conduct simulation studies to study the effect of misspecification of $\Sigma_{*}$ to our procedure. We are interested in how our equicorrelation based procedures react to a different class of error covariance $\Sigma_{*}$, which is assumed to have a corrupted compound symmetry structure that is defined in (25) below. To generate the data, we used the same procedure that is used throughout Section 7 except that we considered different sparsity level of $B_{*}$ (which will be explained below) and used a corrupted compound symmetry $\Sigma_{*}$ which is computed as follows:
$\Sigma_{*}=(1-\omega)\left[0.5\left((1-0.9)I_{q}+0.91_{q}1_{q}^{T}\right)\right]+\omega VDV^{T},$ (25)
where the columns of $V\in\mathbb{R}^{q\times q}$, $v_{1},\ldots,v_{q}$ are generated by the Gram-Schmidt orthogonalization of a $q\times q$ random matrix whose entries are iid standard normal draws; $D\in\mathbb{R}^{q\times q}$ is a diagonal matrix. For the diagonal entries of $D$, we generated iid draws from a two point distribution that is explained below. Through this we vary the condition number of $\Sigma_{*}$. (25) can be interpreted as a convex combination of a general poorly conditioned matrix and a compound symmetry with variance $0.5$ and correlation $0.9$, which is also poorly conditioned. We considered the corruption level $\omega\in\{0.05,0.5,1\}$. $\omega=1$ leads to a general non-sparse $\Sigma_{*}$ that is completely different than compound symmetry. As $\omega$ decreases, the similarity extent of $\Sigma_{*}$ to the compound symmetry class increases. Denoting Ber$(p,a,b)$ by the generic two point distribution that returns a numeric value $a$ (resp. $b$) with the probability $p$ (resp. $1-p$), the following is the condition numbers according to each setting:
  • When $(p,q)=(20,50)$,
    • $D$ constructed upon iid Ber$(0.5,0.1,10)$ draws: 100 / 189.9075 / 412.7517 (setting A / B / C according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid Ber$(0.4,0.01,10)$ draws: 1000 / 507.617 / 453.8222 (setting D / E / F according to $w=1/0.5/0.05$)
  • When $(p,q)=(50,20)$,
    • $D$ constructed upon iid Ber$(0.2,0.2,10)$ draws: 50 / 64.29671 / 153.7895 (setting G / H / I according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid Ber$(0.6,0.02,10)$ draws: 500 / 265.719 / 186.6711 (setting J / K / L according to $w=1/0.5/0.05$)
Note that a $q\times q$ compound symmetry matrix with $q=50$ (resp. $q=20$) and $\theta=0.9$ has condition number 451 (resp. 181). We considered this setting for the elements of $D$ to avoid the diagonally dominant $\Sigma_{*}$, which can be similar to a diagonal matrix. We present numerical experiments that use the diagonally dominant $\Sigma_{*}$ (where the diagonals of $D$ are generated by a chi-square distribution) in H.3 of the supplementary material. For the sparsity levels of $B_{*}$, we considered $(s_{1},s_{2})\in\{(0.5,0.5),\,(0.5,1),\,(1,1)\}$ where the last pair accounts for dense $B_{*}$. For setting A, we denote the sub-setting of sparsity $(0.5,0.5),\,(0.5,1),\,(1,1)$ by A-1, A-2, and A-3, respectively. We use similar notation to denote the sub-settings from B to L. We consider the same set of competitors methods to those described in Section 8.1.
[…]
9 Data analysis
We examined a data set from Korea water resources management information system (WAMIS) on the river flow chart of the Han river and its branches. The entire raw data set is public and can be downloaded from the portal (http://www.wamis.go.kr/wkw/flw_dubobsif.do). We used the daily stream flow measured in $m^{3}/s$ throughout the calendar year 2023 from February 1st to December 25th. The 6 selected measurement points (bridges) for the analysis are: Haengju (HJ), Hankang (HK), Gwangjin (GJ), Paldang (PD), Yeojoo (YJ), Nahmhankang (NHK). The first four points are in the main stream of the Han river, and the next two points are in the relatively northern region. Within these 6 points, when considering the average flux in February to April as a reference, flows in YJ, NHK are generally on the same scale (70–100 $m^{3}/s$), but the other 4 points are generally on much higher scale (120–300 $m^{3}/s)$. We set the predictors as the stream flux observations from 1 to 7 days before and the responses as the daily flux at each measurement point. Hence, there are 42 predictors and the leading intercept term, as well as 6 response variables in the model. We used a training set that has 52 observations from February 8th to March 31th. This indicates $(n,p,q)=(52,43,6)$. The test set has 139 observations from August 8th to December 25th. The competitors are MRCE, ap.MRCE, comb.lasso, sep.lasso, MRCS, MRGCS, ap.MRCS, ap.MRGCS. The labels are defined in Section 7. The prediction performance results are summarized in Table 12. For the grand average prediction taken by all the responses (see Table 1), ap.MRCS was the best model, and MRCS was the second-best. MRCE, ap.MRCE, combined lasso followed the next. ap.MRGCS, MRGCS were slightly outperformed by combined lasso but outperformed separate lasso as well as OLS. ap.MRCS also performed the best in three responses (see Table 2), and MRCS was the best in two responses.
Performance comparison table
Competitors OLS comb.Lasso sep.Lasso MRCS MRGCS ap.MRCS ap.MRGCS MRCE ap.MRCE
River flow 434634.6 265887.3 283362.8 246039.1 268250.1 241223.4 268187.2 261599.1 263000.8
Table 1: The prediction performance comparison table for the combined response. We measured grand averaged squared error from 139 observations in the test set; $\|\hat{Y}-Y\|_{F}^{2}/139$. Boldface indicates the best model [ap.MRCS] and its canonical version [MRCS] which is the second-best.
Performance comparison table by each response
[…]
OLS 1438608 179435.4 217457.6 76625.49 158738.3 536943.0
sep.Lasso 601845.6 246680.3 200497.8 89415.67 147132.8 309751.8
comb.Lasso 601845.6 246680.3 231999.8 95087.78 159095.3 365468.2
MRCS 537848.0 239304.6 200497.8 79507.57 134387.3 284689.6
MRGCS 533929.2 229286.2 210172.7 81730.28 164394.0 389988.1
ap.MRCS 515088.8 233800.8 200497.8 79817.72 134401.6 283733.4
ap.MRGCS 537831.2 224128.6 214104.0 101772.8 164287.5 366998.9
MRCE 525804.0 227598.0 210888.4 80369.34 164309.4 360625.6
ap.MRCE 531651.2 223977.3 215168.6 101128.7 163408.5 342670.8
Table 2: The river flow prediction performance comparison table for each response. We measured average squared error for each response variable from 139 observations in the test set; $(\|\hat{Y}^{i}-Y^{i}\|^{2})/139$, for $1\leq i\leq 6$. Boldface indicates the best model for each response.
10 Acknowledgement
The authors would like to express gratitude to Professor Aaron Molstad for helpful comments on the proof details. Computational resources were provided by the WVU Research Computing Thorny Flat HPC cluster, which is funded in part by NSF OAC-1726534. Price was partially supported by the National Institute of General Medical Sciences, 5U54GM104942-04. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix A Derivation of the optimization problems (4)
We first check the derivation of the optimization problem (4). Recall that
$\Sigma_{*}=\eta^{2}_{*}\left\{(1-\theta_{*})I_{q}+\theta_{*}1_{q}1_{q}^{T}\right\}$
Since ${\rm det}(I+uv^{T})=1+u^{T}v$ for $u,\,v\in\mathbb{R}^{q}$ and ${\rm det}(cA)=c^{q}A$ for $A\in\mathbb{R}^{q\times q}$, we have
${\rm det}(\Sigma_{*})$ $={\rm det}\left(\eta^{2}_{*}(1-\theta_{*})\left[I_{q}+\frac{\theta_{*}}{1-\theta_{*}}1_{q}1_{q}^{T}\right]\right)$
$=\eta_{*}^{2q}(1-\theta_{*})^{q}\left(1+q\frac{\theta_{*}}{1-\theta_{*}}\right).$
This yields
$\log{\rm det}(\Sigma_{*})$ $=q\log(\eta_{*}^{2})+q\log(1-\theta_{*})+\log\left(1+q\frac{\theta_{*}}{1-\theta_{*}}\right)$
$=q\log(\eta_{*}^{2})+q\log(1-\theta_{*})+\log\left(\frac{1-\theta_{*}+q\theta_{*}}{1-\theta_{*}}\right)$
$=q\log(\eta_{*}^{2})+(q-1)\log(1-\theta_{*})+\log(1+(q-1)\theta_{*}).$
To compute $\Sigma_{*}^{-1}$, we use the following so-called Woodbury identity:
$(A+CDC^{\prime})^{-1}=A^{-1}-A^{-1}C(D^{-1}+C^{\prime}A^{-1}C)^{-1}C^{\prime}A^{-1},$
with $A=(1-\theta_{*})I_{q}$, $D=\theta_{*}$, and $C=1_{q}$. Then $A^{-1}=(1-\theta_{*})^{-1}I_{q}$, $D^{-1}=\theta_{*}^{-1}$, and
$\Sigma_{*}^{-1}$ $=(\eta_{*}^{2})^{-1}\left((1-\theta_{*})I_{q}+\theta_{*}1_{q}1_{q}^{T}\right)^{-1}$
$=(\eta_{*}^{2})^{-1}(A+CDC^{T})^{-1}$
$=(\eta_{*}^{2})^{-1}\left[A^{-1}-A^{-1}C(D^{-1}+C^{\prime}A^{-1}C)^{-1}C^{T}A^{-1}\right]$
$=(\eta_{*}^{2})^{-1}\left[(1-\theta_{*})^{-1}I_{q}-(1-\theta_{*})^{-1}I_{q}1_{q}(\theta_{*}^{-1}+(1-\theta_{*})^{-1}1_{q}^{T}I_{q}1_{q})^{-1}1_{q}^{T}(1-\theta_{*})^{-1}I_{q}\right]$
$=(\eta_{*}^{2})^{-1}\left[(1-\theta_{*})^{-1}I_{q}-(1-\theta_{*})^{-1}I_{q}1_{q}(\theta_{*}^{-1}+(1-\theta_{*})^{-1}q)^{-1}1_{q}^{T}(1-\theta)^{-1}I_{q}\right]$
$=(\eta_{*}^{2})^{-1}\left[(1-\theta_{*})^{-1}I_{q}-(1-\theta_{*})^{-2}(\theta^{-1}_{*}+(1-\theta_{*})^{-1}q)^{-1}1_{q}1_{q}^{T}\right]$
$=(\eta_{*}^{2})^{-1}(1-\theta_{*})^{-1}\left(I_{q}-\frac{\theta_{*}}{1-\theta_{*}+q\theta_{*}}1_{q}1_{q}^{T}\right).$
This identifies
$\begin{split}F_{\lambda}(B,\eta^{2},\theta;Y,X)&=\frac{1}{\eta^{2}(1-\theta)}\operatorname{tr}\left\{n^{-1}(Y-XB)^{T}(Y-XB)\left[I_{q}-\frac{\theta}{1-\theta+q\theta}1_{q}1^{T}_{q}\right]\right\}\\ &\indent+(q-1)\log(1-\theta)+\log(1+\{q-1\}\theta)+q\log(\eta^{2})\\ &\indent+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|.\end{split}$ (26)
Its equivalence to the optimization problem (4) is based on the following:
$\begin{split}&\frac{1}{\eta^{2}(1-\theta)}\operatorname{tr}\left\{n^{-1}(Y-XB)^{T}(Y-XB)\left[I_{q}-\frac{\theta}{1-\theta+q\theta}1_{q}1^{T}_{q}\right]\right\}\\ &=\frac{1}{n\eta^{2}(1-\theta)}\|Y-XB\|_{F}^{2}-\frac{\theta}{n\eta^{2}(1-\theta)(1-\theta+q\theta)}\operatorname{tr}\left\{(Y-XB)^{T}(Y-XB)1_{q}1_{q}^{T}\right\}\\ &=\frac{1}{n\eta^{2}(1-\theta)}\|Y-XB\|_{F}^{2}-\frac{\theta}{n\eta^{2}(1-\theta)(1-\theta+q\theta)}\|(Y-XB)1_{q}\|^{2}.\end{split}$
We now check the derivation of the optimization problem (11). We cyclically update $\eta_{j}$ while fixing $B$, $\theta$, and $\eta_{-j}$, where $\eta_{-j}=(\eta_{1},\ldots,\eta_{j-1},\eta_{j+1},\ldots,\eta_{q})$ for each $j\in\{1,\ldots,q\}$. By simple algebra, one can derive that the objective function with respect to $\eta_{j}$ is computed as the following:
$\begin{split}g_{j}(\eta_{j})&=\frac{1}{n(1-\theta)}\sum_{i=1}^{n}e_{ij}^{2}\frac{1}{\eta_{j}^{2}}-\frac{\theta}{n(1-\theta)(1+(q-1)\theta)}\sum_{i=1}^{n}\left(\sum_{j=1}^{q}e_{ij}\frac{1}{\eta_{j}}\right)^{2}+2\log(\eta_{j}),\end{split}$
We suppress the dependence of $g_{j}$ on $\eta_{-j}$, $B$, and $\theta$ for notational simplicity. From the above computation, we have
$\frac{1}{2}g_{j}^{\prime}(\eta_{j})$ $=\frac{1}{\eta_{j}}+\frac{\theta}{n(1-\theta)(1+(q-1)\theta)}\sum_{i=1}^{n}\left(\sum_{k\neq j}e_{ij}e_{ik}\frac{1}{\eta_{j}^{2}\eta_{k}}\right)$
$\indent-\frac{1+(q-2)\theta}{n(1+(q-1)\theta)(1-\theta)}\left(\sum_{i=1}^{n}e_{ij}^{2}\right)\frac{1}{\eta_{j}^{3}}.$
Then $g_{j}^{\prime}=0$ yields (12), since $\eta_{j}>0$.
Appendix B Proof of Theorem 1
We first provide two lemmas which are used for the proof of Theorem 1. Recall $S=\{i:{\rm vec}(B_{*})_{i}\neq 0\}$, and
$\Omega_{*}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}\left[I_{q}-\frac{\theta_{*}}{1+(q-1)\theta_{*}}1_{q}1_{q}^{T}\right].$
Lemma 1.
Under the conditions (A1) and (A3), assuming that $(\hat{\eta}^{2},\hat{\theta})$ are $\sqrt{n}$-consistent estimators of $(\eta_{*}^{2},\theta_{*})$, then, if $\lambda\sqrt{n}\rightarrow 0$ and $\lambda n^{(r+1)/2}\rightarrow\infty$, the following holds.
$\lim_{n\rightarrow\infty}P(\hat{B}_{jk}=0)=1\qquad\qquad{\rm if~{}}B_{jk*}=0,$
$\sqrt{n}({\rm vec}(\hat{B})_{S}-{\rm vec}(B_{*})_{S})\rightarrow_{d}N(0,D^{-1}),$
where $\hat{B}$ is a solution to the optimization problem (14) for fixed $(\hat{\eta}^{2},\hat{\theta})$, and $D=(\Omega_{*}\otimes Z)_{S}$.
Proof.
The diagonals of $\Omega_{*}$ are $\frac{1+(q-2)\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}(1-\frac{\theta_{*}}{1+(q-1)\theta_{*}})$ and the off-diagonals are $-\frac{\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}$. Let $v_{*}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}$ and $w_{*}=\frac{\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}$. We further denote $\hat{v}=\frac{1}{\hat{\eta}^{2}(1-\hat{\theta})}$ and $\hat{w}=\frac{\hat{\theta}}{\hat{\eta}^{2}(1-\hat{\theta})(1+(q-1)\hat{\theta})}$. Then,
$|\hat{v}-v_{*}|$ $=\left|\frac{\hat{\eta}^{2}(1-\hat{\theta})-\eta^{2}_{*}(1-\theta_{*})}{\eta^{2}_{*}(1-\theta_{*})\hat{\eta}^{2}(1-\hat{\theta})}\right|$
$\leq\frac{|\hat{\eta}^{2}-\eta^{2}_{*}|+\hat{\eta}^{2}|\hat{\theta}-\theta_{*}|+\theta_{*}|\hat{\eta}^{2}-\eta^{2}_{*}|}{\eta^{2}_{*}(1-\theta_{*})\hat{\eta}^{2}(1-\hat{\theta})}=O_{p}(1/\sqrt{n}),$
since $\hat{\eta}^{2},1/\hat{\eta}^{2},1/(1-\hat{\theta})=O_{p}(1)$ and $|\hat{\eta}^{2}-\eta^{2}_{*}|,|\hat{\theta}-\theta_{*}|=O_{p}(1/\sqrt{n})$. Similarly, one can check that $|\hat{w}-w_{*}|=O_{p}(1/\sqrt{n})$. This implies that the $\sqrt{n}$-consistent estimator of $(\eta^{2}_{*},\theta_{*})$ guarantees the existence of $\sqrt{n}$-consistent estimator of $\Omega$. The rest of the proof follows the same derivation steps provided in the proofs of Lemma 1 and Theorem 1 of Lee and Liu [6]. Hence, we omit the proof. ∎
We define a matrix, $V\in\mathbb{R}^{2\times 2}$, which characterizes the limiting covariance of $(\hat{\eta},\hat{\theta})$. This symmetric matrix has the following four elements:
$\begin{split}&V_{11}=\frac{1}{q^{2}}\left(\sum_{j=1}^{q}{\rm var}(E_{ij}^{2})+2\sum_{j<k}{\rm cov}(E_{ij}^{2},E_{ik}^{2})\right),\\ &V_{22}=\frac{1}{q^{2}(q-1)^{2}}\sum_{j,k,l,m}Q_{jk}Q_{lm}{\rm cov}(E_{ij}E_{ik},E_{il}E_{im}),\\ &V_{12}=V_{21}=\frac{1}{q^{2}}{\rm var}\left(\sum_{j=1}^{q}E_{ij}^{2}\right)-\frac{2}{q^{2}(q-1)}{\rm cov}\left(\sum_{j<k}E_{ij}E_{ik},\sum_{j=1}^{q}E_{ij}^{2}\right),\end{split}$ (27)
where $Q=(qI_{q}-1_{q}1_{q}^{T})$.
Lemma 2.
Under the conditions (A1), (A2) and (A4), assuming that $\hat{B}$ is $\sqrt{n}$-consistent to $B_{*}$, then, for $W=(1,-1/\eta_{*}^{2})^{T}$, the following holds.
$\left\|(\hat{\eta}^{2},\hat{\theta})-(\eta_{*}^{2},\theta_{*})\right\|=O_{p}(1/\sqrt{n}),$
where $(\hat{\eta}^{2},\hat{\theta})$ is a solution to the optimization problem (14) for fixed $\hat{B}$. Furthermore, if $\theta_{*}\in(0,1)$, then
$\sqrt{n}(\hat{\eta}^{2},\hat{\theta})-(\eta_{*}^{2},\theta_{*}))\rightarrow_{d}W^{T}N_{2}(0,V),$
where $W=(1,-1/\eta_{*}^{2})^{T}$, and $V$ is defined in 27.
Proof.
It suffices to assume that true $B_{*}$ is known, since the natural following step which is the replacement of $B_{*}$ with $\hat{B}$ can be directly obtained by application of the derivation used in the proof of Theorem 2 in Lee and Liu [6]. Recall the reparametrization of $(\eta^{2},\theta)$ to $(\alpha,\gamma)$ via $\alpha=\eta^{2}(1-\theta)$, $\gamma=\eta^{2}(1+(q-1)\theta)$ by which we derived the closed form solutions to $\hat{\alpha}$ and $\hat{\gamma}$. The solutions were $\hat{\alpha}=\frac{qM_{1}^{*}-M_{2}^{*}}{q(q-1)}$, $\hat{\gamma}=\max\{\hat{\alpha},M_{2}^{*}/q\}$, where $M_{1}^{*}=\frac{1}{n}\|E\|_{F}^{2}$ and $M_{2}^{*}=\frac{1}{n}\|E1_{q}\|^{2}$ (see (6), (7)). We define $\hat{\delta}=\frac{M_{1}^{*}}{q}$. We first find the joint limiting distribution of $(\hat{\delta},\hat{\alpha})$. We have
$\frac{qM_{1}^{*}-M_{2}^{*}}{q(q-1)}$ $=\frac{1}{nq(q-1)}\left\{\sum_{i=1}^{n}\left[\sum_{j=1}^{q}(q-1)E_{ij}^{2}-2\sum_{1\leq j<k\leq q}E_{ij}E_{ik}\right]\right\}$
$=\frac{1}{nq(q-1)}\sum_{i=1}^{n}E_{i}^{T}QE_{i},$
where $E_{i}=(E_{i1},\ldots,E_{iq})^{T}\in\mathbb{R}^{q}$ and $Q=(qI_{q}-1_{q}1_{q}^{T})$. We further have
$\mathbb{E}(E_{i}^{T}QE_{i}/(q(q-1)))=\eta_{*}^{2}(1-\theta_{*}),$
${\rm var}\left(E_{i}^{T}QE_{i}/(q(q-1))\right)=\frac{1}{q^{2}(q-1)^{2}}\sum_{j,k,l,m}Q_{jk}Q_{lm}{\rm cov}(E_{ij}E_{ik},E_{il}E_{im}).$
In addition, $\hat{\delta}$ can be expressed as $\frac{1}{nq}\sum_{i=1}^{n}\sum_{j=1}^{q}E_{ij}^{2}=\frac{1}{nq}\sum_{i=1}^{n}E_{i}^{T}I_{q}E_{i}$, which satisfies
$\mathbb{E}(E_{i}^{T}I_{q}E_{i}/q)=\eta_{*}^{2},$
${\rm var}\left(E_{i}^{T}I_{q}E_{i}/q\right)=\frac{1}{q^{2}}\left(\sum_{j=1}^{q}{\rm var}(E_{ij}^{2})+2\sum_{j<k}{\rm cov}(E_{ij}^{2},E_{ik}^{2})\right).$
Likewise, we obtain
${\rm cov}\left(E_{i}^{T}I_{q}E_{i}/q,E_{i}^{T}QE_{i}/(q(q-1))\right)=\frac{1}{q^{2}}{\rm var}\left(\sum_{j=1}^{q}E_{ij}^{2}\right)-\frac{2}{q^{2}(q-1)^{2}}{\rm cov}\left(\sum_{j<k}E_{ij}E_{ik},\sum_{j=1}^{q}E_{ij}^{2}\right).$
Thus, we have
$\sqrt{n}((\hat{\delta},\hat{\alpha})-(\eta_{*}^{2},\eta_{*}^{2}(1-\theta_{*})))\rightarrow_{d}N_{2}(0,V).$ (28)
And, since $\hat{\eta}^{2}=(1-1/q)\hat{\alpha}+\hat{\gamma}/q$, we have
$\hat{\eta}^{2}-\hat{\delta}=\frac{\hat{\gamma}}{q}-\frac{M_{2}^{*}}{q^{2}}=\max\left\{\frac{\hat{\alpha}}{q},\frac{M_{2}^{*}}{q^{2}}\right\}-\frac{M_{2}^{*}}{q^{2}}.$
When $\theta_{*}\in(0,1)$, $\hat{\alpha}\rightarrow_{p}\eta^{2}(1-\theta)$, and $\max\{\hat{\alpha},M_{2}^{*}/q\}=M_{2}^{*}/q$ with probability tending to $1$, since $M_{2}^{*}/q\rightarrow_{p}\eta_{*}^{2}$. This yields, $\hat{\eta}^{2}=\hat{\delta}$ with probability tending to $1$ and it further implies
$\sqrt{n}((\hat{\eta}^{2},\hat{\alpha})-(\eta^{2}_{*},\eta^{2}_{*}(1-\theta_{*})))\rightarrow_{d}N_{2}(0,V).$
Then, by the same Delta method via $g(x,y)=(x,(x-y)/x)$, we have
$\sqrt{n}\left(\left(\hat{\eta}^{2},\frac{\hat{\eta}^{2}-\hat{\alpha}}{\hat{\eta}^{2}}\right)-(\eta_{*}^{2},\theta_{*})\right)\rightarrow_{d}W^{T}N_{2}(0,V),$
where $W=(1,-1/\eta_{*}^{2})^{T}$, since $\hat{\delta}>0$ almost surely. It is obvious that $\frac{\hat{\eta}^{2}-\hat{\alpha}}{\hat{\eta}^{2}}=\frac{\hat{\gamma}-\hat{\alpha}}{\hat{\gamma}+(q-1)\hat{\alpha}}=\hat{\theta}$. On the other hand, when $\theta_{*}=0$, (28) with the Delta method yields,
$\sqrt{n}\left(\left(\hat{\delta},\frac{\hat{\delta}-\hat{\alpha}}{\hat{\delta}}\right)-(\eta_{*}^{2},\theta_{*})\right)\rightarrow_{d}W^{T}N_{2}(0,V).$
Then it suffices to show $|\hat{\delta}-\hat{\eta}^{2}|=O_{p}(1/\sqrt{n})$ and $|\hat{\alpha}(1/\hat{\eta}^{2}-1/\hat{\delta})|=O_{p}(1/\sqrt{n})$. Indeed, we have
$q|\hat{\delta}-\hat{\eta}^{2}|\leq\left|\frac{M_{2}^{*}}{q}-\hat{\alpha}\right|=\left|\frac{M_{2}^{*}}{q}-\frac{qM_{1}^{*}-M_{2}^{*}}{q(q-1)}\right|=\left|\frac{M_{2}^{*}-M_{1}^{*}}{q-1}\right|,$
and $M_{2}^{*}-M_{1}^{*}=\frac{1}{n}\sum_{i=1}^{n}E_{i}^{T}Q_{0}E_{i}$, where $Q_{0}$ has 0 for all diagonal elements and $1$ for all off-diagonal elements. Thus, by similar asymptotic derivation steps above, one can easily check $|\hat{\delta}-\hat{\eta}^{2}|=O_{p}(1/\sqrt{n})$. Since $\hat{\eta}^{2},\hat{\delta},\hat{\alpha}\rightarrow_{p}\eta_{*}^{2}>0$, $|1/\hat{\eta}^{2}|,|1/\hat{\delta}|,|\hat{\alpha}|=O_{p}(1)$. This implies
$|\hat{\alpha}(1/\hat{\eta}^{2}-1/\hat{\delta})|=|\hat{\alpha}||1/\hat{\delta}||1/\hat{\eta}^{2}||\hat{\eta}^{2}-\hat{\delta}|=O_{p}(1/\sqrt{n}).$
Finally, we have
$\left\|\left(\hat{\delta},\frac{\hat{\delta}-\hat{\alpha}}{\hat{\delta}}\right)-\left(\hat{\eta}^{2},\frac{\hat{\eta}^{2}-\hat{\alpha}}{\hat{\eta}^{2}}\right)\right\|\leq|\hat{\eta}^{2}-\hat{\delta}|+|\hat{\alpha}(1/\hat{\eta}^{2}-1/\hat{\delta})|=O_{p}(1/\sqrt{n}),$
and it completes the proof. ∎
Now we give the proof of Theorem 1.
Proof.
We already verified that a $\sqrt{n}$-consistent estimator of $(\eta_{*}^{2},\theta_{*})$ guarantees the existence of the $\sqrt{n}$-consistent estimator of $\Omega_{*}$ by Lemma 1. Combining the result in Lemma 2, the rest of the proof follows the same derivation used in the proofs of Lemma 3 and Theorem 3 of Lee and Liu [6]. Hence, we omit the proof. ∎
Appendix C High-dimensional statistical guarantees
In this section, we study the statistical properties of an oracle estimator, MRCS-Or that assumes the true covariance is known (see (19)), and a surrogate version of ap.MRCS (see Section 3.2) that leverages pilot estimators $\eta^{2}_{n}$ and $\theta_{n}$ for $\eta^{2}_{*}$ and $\theta_{*}$. Our theoretical results accommodate scenarios where the sample size $n$ and the dimensional parameters $p$ and $q$ are allowed to grow simultaneously.
For a matrix $M\in\mathbb{R}^{p\times q}$, we let ${\rm vec}(M)\in\mathbb{R}^{pq}$ denote the vector formed by stacking the columns of $M$. We denote $i$-th row of $M$ by $M_{i\cdot}$, and $j$-th column of it by $M_{\cdot j}$. Recall $S=\{i:{\rm vec}(B_{*})_{i}\neq 0\}$, which is the support of the true regression coefficients matrix $B_{*}$. We define the subspace that corresponds to $S$ as follows:
$\mathcal{M}(S)\equiv\{v\in\mathbb{R}^{pq}:v_{i}=0~{}{\rm if}~{}i\notin S\}.$
We further define a projection of a generic verctor $u$ onto the subspace $\mathcal{M}(S)$ by
$u_{M(S)}\equiv\mathop{arg\,min}\displaylimits_{v\in\mathcal{M}(S)}\|u-v\|_{2}.$
We let $\mathcal{M}^{\perp}(S)$ denote the orthogonal complement of $\mathcal{M}(S)$, which is given by
$\mathcal{M}^{\perp}(S)\equiv\{v^{pq}:\langle u,v\rangle=0,\,\forall u\in\mathcal{M}(S)\}.$
To prove consistency of the oracle procedure, we make the following assumptions:
  1. (C1)
    : Each column of $X$ is scaled so that $\|X_{\cdot j}\|_{2}=\sqrt{n}$ for all $j=1,\ldots,p$.
  2. (C2)
    : The errors $E_{i}=(E_{i1},\ldots,E_{iq})$ are iid $N_{q}(0,\Sigma_{*})$, for $i=1,\ldots,n$, where
    $\Sigma_{*}=\eta^{2}_{*}\left[(1-\theta_{*})I_{q}+\theta_{*}1_{q}1_{q}^{T}\right],$
    for $\eta_{*}\in(0,\infty)$ and $\theta_{*}\in[0,1)$.
  3. (C3)
    : Let $\bar{X}=I_{q}\otimes X$, where $\otimes$ stands for the standard Kronecker product. There exists a positive constant $\kappa$ which satisfies
    $\kappa\|v\|_{2}^{2}\leq\min_{v\in\mathcal{C}}\|\bar{X}v\|_{2}^{2}/n,$
    where
    $\mathcal{C}:=\{v\in\mathbb{R}^{pq}|\|v_{\mathcal{M}^{\perp}(S)}\|_{1}\leq 3\|v_{\mathcal{M}}\|_{1}\}.$ (29)
Assumption (C1) is a standard assumption in the penalized regression literature [8, 9, 15]. This condition can be easily satisfied with appropriate standardization. Assumption (C2) implies that the error follows multivariate normal distribution, which has zero mean and exhibits compound symmetry variance. Assumption (C3) is so-called the restricted eigenvalue condition. We refer the readers to Negahban et al. [8] for more details of this condition. Now we state the consistency result of $\hat{B}_{Or}$.
Theorem 2.
If Assumptions (C1)(C3) hold and $\lambda_{n}=8\sqrt{\frac{\log(pq)}{n\eta^{2}_{*}(1-\theta_{*})}}$, then there exist positive constants $c_{1},\,c_{2},\,c_{3}$ such that
$\|{\rm vec}(\hat{B}_{\rm Or}-B*)\|_{2}\leq c_{3}\frac{\eta_{*}}{\sqrt{1-\theta_{*}}}\frac{1+(q-1)\theta_{*}}{\kappa}\sqrt{\frac{s\log(pq)}{n}},$
with probability at least $1-c_{1}\exp(-c_{2}\log(pq))$.
Proof of Theorem 2 is postponed to D of the supplementary material. Except for the scalar factor $(1+(q-1)\theta_{*})/\sqrt{1-\theta_{*}}$, which is related to the eigenspectrum of the correlation matrix of $\Sigma_{*}$, the rate $O_{p}(\sqrt{s\log(pq)/n})$ is standard rate of convergence in lasso-penalized multivariate linear regression [9, 17, 13]. However, since the largest eigenvalue of $\Sigma_{*}$ is proportional to $q$, it is required that $q\sqrt{s\log(pq)/n}\rightarrow 0$ when $\theta_{*}>0$ to obtain $\|{\rm vec}(\hat{B}_{\rm Or}-B*)\|_{2}=o_{p}(1)$.
Assume that we have consistent pilot estimators $\eta^{2}_{n},\,\theta_{n}$ of $\eta^{2}_{*},\,\theta_{*}$. We consider the following updated optimization problem of (19) with replacement of $\eta^{2}_{*},\,\theta_{*}$ with plug-in pilot estimators $\eta^{2}_{n},\,\theta_{n}$.
$\begin{split}\hat{B}_{Pl}:=\mathop{arg\,min}\displaylimits_{B\in\mathbb{R}^{p\times q}}&\operatorname{tr}\left\{n^{-1}(Y-XB)^{T}(Y-XB)\left[\eta^{2}_{n}\left\{\theta_{n}1_{q}1_{q}^{T}+(1-\theta_{n})I_{q}\right\}\right]^{-1}\right\}\\ &\indent+\lambda\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|\end{split}$ (30)
We define
$v(\eta^{2},\theta):=\frac{1}{\eta^{2}(1+(q-1)\theta)},$ (31)
for $\eta^{2}\in(0,\infty)$ and $\theta\in[0,1)$. To prove consistency of the plug-in optimizer, we make the following assumption on $\eta^{2}_{n},\,\theta_{n}$.
  1. (C4)
    : There exists pilot estimators $\eta^{2}_{n},\,\theta_{n}$ for $\eta^{2}_{*},\,\theta_{*}$ that are independent of $E_{1},\ldots,E_{n}$. The pilot estimators satisfy $\eta_{n}\in(0,\infty)$, $\theta_{n}\in[0,1)$, and the following holds with probability at least $1-k_{1}\exp(-k_{2}\log(pq))$ where $k_{1}>0,\,k_{2}\geq 1$:
    $\left|\frac{v(\eta^{2}_{n},\theta_{n})}{v(\eta^{2}_{*},\theta_{*})}-1\right|\leq\frac{1}{2},$ (32)
    and
    $\begin{split}s_{n}:=\left[(1-\theta_{*})+|\theta_{n}-\theta_{*}|\right]\frac{(\eta^{2}_{*})^{3}}{(\eta^{2}_{n})^{2}}\frac{(1-\theta_{*})^{2}}{(1-\theta_{n})^{2}}\leq 2\eta^{2}_{*}(1-\theta_{*}).\end{split}$ (33)
Assumption (C4) holds if $\eta^{2}_{n},\,\theta_{n}$ are sufficiently close to $\eta^{2}_{*},\,\theta_{*}$, respectively, with high probability. The independence between $(\eta^{2}_{n},\,\theta_{n})$ and $E_{1},\ldots,E_{n}$ can be achieved when $(\eta^{2}_{n},\,\theta_{n})$ are estimated from independent observations generated by the same model. More details can be found in E of the supplementary material.
While the following theorem is not directly applicable to ap.MRCS (see Section 3.2), the theoretical properties of $\hat{B}_{Pl}$ provide better understanding of ap.MRCS. Now we state the consistency result of $\hat{B}_{Pl}$.
Theorem 3.
If Assumptions (C1)(C4) hold and $\lambda_{n}=8\sqrt{\frac{\log(pq)}{n\eta^{2}_{*}(1-\theta_{*})}}$, then there exist constants $c_{4},\,c_{5},\,c_{6}$ such that
$\|{\rm vec}(\hat{B}_{Pl}-B_{*})\|_{2}\leq c_{6}\frac{\eta_{*}}{\sqrt{1-\theta_{*}}}\frac{1+(q-1)\theta_{*}}{\kappa}\sqrt{\frac{s\log(pq)}{n}},$
with probability at least $1-c_{4}\exp(-c_{5}\log(pq))$.
Proof of Theorem 3 is provided in F of the supplementary material. The convergence rate of $\|{\rm vec}(\hat{B}_{Pl}-B_{*})\|_{2}$ is identical to the rate of $\|{\rm vec}(\hat{B}_{Or}-B_{*})\|_{2}$. On the other hand, compared with the upper bound of $\|{\rm vec}(\hat{B}_{Or}-B_{*})\|_{2}$ in Theorem 2, the upper bound of $\|{\rm vec}(\hat{B}_{Pl}-B_{*})\|_{2}$ in Theorem 3 is inflated by $\sqrt{2}$ (see the proof given in F), which arises from (33) that accounts for additional error introduced by plug-in pilot estimators.
Appendix D Proof of Theorem 2
In this section, we give the proof of Theorem 2.
Proof.
Let $\bar{Y}={\rm vec}(Y)$, $\bar{B}={\rm vec}(B)$, and $\bar{E}={\rm vec}(E)$. Letting $\Omega_{*}=\Sigma_{*}^{-1}$, we can write (19) as
$\hat{B}_{\rm Or}:=\mathop{arg\,min}\displaylimits\mathcal{L}(B)+\lambda_{n}\|\bar{B}\|_{1},$ (34)
where
$\mathcal{L}(B)=\frac{1}{n}(\bar{Y}-\bar{X}\bar{B})^{T}(\Omega_{*}\otimes I_{n})(\bar{Y}-\bar{X}\bar{B}).$
We define $\delta\mathcal{L}(\Delta,\bar{B})=\mathcal{L}(\bar{B}+\Delta)-\mathcal{L}(\bar{B})-\langle\nabla\mathcal{L}(\bar{B}),\Delta\rangle$. Since $\Omega_{*}\otimes I_{n}$ is positive definite symmetric matrix with minimal eigenvalue of $1/(\eta^{2}_{*}(1+(q-1)\theta_{*}))$, one can easily show
$\delta\mathcal{L}(\Delta,\bar{B})=\frac{1}{n}(\bar{X}\Delta)^{T}(\Omega_{*}\otimes I_{n})(\bar{X}\Delta)$ $\geq\frac{1}{(1+(q-1)\theta_{*})\eta^{2}_{*}}\frac{\|\bar{X}\Delta\|_{2}^{2}}{n}$ (35)
$\geq\frac{\kappa\|\Delta\|_{2}^{2}}{(1+(q-1)\theta_{*})\eta^{2}_{*}},$ (36)
for all $\Delta\in\mathcal{C}$, by Assumption (C3). For a general subspace $\mathcal{X}$, we define a compatibility constant $\psi(\mathcal{M})=\sup_{u\in\mathcal{M}\backslash\{0\}}\frac{\|u\|_{1}}{\|u\|_{2}}$, where $\psi(\mathcal{M}(S))=\sqrt{s}$. If $\lambda_{n}\geq 2\|\nabla\mathcal{L}(\bar{B})\|_{\infty}$, by Theorem 1 of Negahban et al. [8], we have
$\|{\rm vec}(\hat{B}_{\rm Or}-B_{*})\|_{2}^{2}\leq 9\frac{\lambda_{n}^{2}}{c^{2}}s,$ (37)
where $c=\kappa/((1+(q-1)\theta_{*})\eta_{*}^{2})$. Indeed,
$\|\nabla\mathcal{L}(\bar{B})\|_{\infty}$ $=\frac{2}{n}\|\bar{X}^{T}(\Omega_{*}\otimes I_{n})(\bar{Y}-\bar{X}\bar{B})\|_{\infty}$
$=\frac{2}{n}\|\bar{X}^{T}(\Omega_{*}\otimes I_{n})\bar{E}\|_{\infty}=\frac{2}{n}\|\bar{X}^{T}\bar{W}\|_{\infty},$
where $(\Omega_{*}\otimes I_{n})\bar{E}=\bar{W}$; $\bar{W}:={\rm vec}(W)$, $W\in\mathbb{R}^{n\times q}$, and $W_{i\cdot}$’s are iid from $N(0,\Omega_{*})$, for $i=1,\ldots,n$. We define
$u(\eta^{2},\theta):=\frac{1}{\eta^{2}(1-\theta)}\frac{1+(q-2)\theta}{1+(q-1)\theta},$ (38)
for $\eta\in(0,\infty)$ and $\theta\in[0,1)$. We define the events
$\mathcal{A}_{j}=\left\{\frac{1}{n}\|X^{T}W_{\cdot j}\|_{\infty}\leq\sqrt{u(\eta^{2}_{*},\theta_{*})\frac{t+2\log p}{n}}\right\},$
$\mathcal{A}=\cap_{j=1}^{q}\mathcal{A}_{j},.$
Since $\mathbb{P}(\mathcal{A}_{j}^{c})\leq 2\exp(-t/2)$ by Lemma 6.2 of Bühlmann and Van De Geer [2] and Assumption (C1)(C2), we have
$\mathbb{P}(\mathcal{A})=\mathbb{P}(\cap_{j=1}^{q}\mathcal{A}_{j})\geq 1-\sum_{j=1}^{q}\mathbb{P}(\mathcal{A}_{j}^{c})\geq 1-2q\exp(-t/2).$
Recall that $\lambda_{n}=8\sqrt{\frac{\log(pq)}{n\eta^{2}_{*}(1-\theta_{*})}}$. Thus, letting $t=4\log q+2\log p$, we have
$\begin{split}\mathbb{P}\left(\lambda_{n}\geq 2\|\nabla\mathcal{L}(\bar{B})\|_{\infty}\right)&=\mathbb{P}\left(\lambda_{n}\geq\frac{4}{n}\|\bar{X}^{T}\bar{W}\|_{\infty}\right)\\ &\geq 1-2\exp(-\log(pq)),\end{split}$ (39)
since $\frac{1+(q-2)\theta_{*}}{1+(q-1)\theta_{*}}\leq 1$. Combining (37) and (39) completes the proof with $c_{1}=2,\,c_{2}=1,\,c_{3}=24$. ∎
Appendix E Supplementary description of Assumption (C4)
First, we note that
$\left|\frac{v(\eta^{2}_{*},\theta_{*})}{v(\eta^{2}_{n},\theta_{n})}-1\right|$
$\indent=\left|\frac{(1+(q-1)\theta_{*})(\eta^{2}_{n}-\eta^{2}_{*})+(q-1)\eta^{2}_{n}(\theta_{n}-\theta_{*})}{\eta^{2}_{*}(1+(q-1)\theta_{*})}\right|$
$\indent=\left|\frac{\eta^{2}_{n}}{\eta^{2}_{*}}-1\right|+q\frac{\eta^{2}_{n}}{\eta^{2}_{*}}\frac{|\theta_{n}-\theta_{*}|}{1+(q-1)\theta_{*}}$
$\indent\leq\left|\frac{\eta^{2}_{n}}{\eta^{2}_{*}}-1\right|+\left(\left|\frac{\eta^{2}_{n}}{\eta^{2}_{*}}-1\right|+1\right)\frac{q|\theta_{n}-\theta_{*}|}{1+(q-1)\theta_{*}}.$
Thus, if
$\left|\eta^{2}_{n}-\eta^{2}_{*}\right|\leq\frac{\eta^{2}_{*}}{6},~{}{\rm and}\quad\left|\theta_{n}-\theta_{*}\right|\leq\frac{1+(q-1)\theta_{*}}{7q}$ (40)
on an event $\mathcal{A}_{0}$ where $P(\mathcal{A}_{0})\geq 1-k_{1}\exp{(-k_{2}\log(pq))}$ for $k_{1}>0$ and $k_{2}\geq 1$, then $\left|\frac{v(\eta^{2}_{n},\theta_{n})}{v(\eta^{2}_{*},\theta_{*})}-1\right|\leq 1/3$ on $\mathcal{A}_{0}$. This further yields $\left|\frac{v(\eta^{2}_{*},\theta_{*})}{v(\eta^{2}_{n},\theta_{n})}-1\right|\leq 1/2$ on $\mathcal{A}_{0}$, which is (32).
Next, when
$\frac{\eta^{2}_{*}}{\eta^{2}_{n}}\leq\left(\frac{3}{2}\right)^{1/4},\quad\frac{1-\theta_{*}}{1-\theta_{n}}\leq\left(\frac{3}{2}\right)^{1/4},{\rm~{}and}\quad|\theta_{n}-\theta_{*}|\leq\frac{1-\theta_{*}}{3}$ (41)
hold on $\mathcal{A}_{0}$, one can easily check that (33) holds on $\mathcal{A}_{0}$. Thus, when $\mathbb{E}|\eta^{2}_{n}-\eta^{2}_{*}|=O(1/(pq)^{k_{2}})$ and $\mathbb{E}|\theta_{n}-\theta_{*}|=O(q/(pq)^{k_{2}})$, then the conditions in (40)–(41) hold.
Appendix F Proof of Theorem 3
In this section, we give the proof of Theorem 3.
Proof.
As in the proof of Theorem 2, we denote $\bar{Y}={\rm vec}(Y)$, $\bar{B}={\rm vec}(\hat{B})$, and $\bar{E}={\rm vec}(E)$. We consider the following derivation conditioned on an event $\mathcal{C}_{0}$ that satisfies (32)–(33), which further satisfies $P(\mathcal{C}_{0})\geq 1-k_{1}\exp(-k_{2}\log(pq))$. And, we write
$\begin{split}\hat{B}_{Pl}:=\mathop{arg\,min}\displaylimits_{B\in\mathbb{R}^{p\times q}}\mathcal{L}_{n}(B)+\lambda_{n}\sum_{j=1}^{p}\sum_{k=1}^{q}|B_{jk}|\end{split}$ (42)
Since $\hat{B}_{Pl}$ is the minimizer of the problem (42), we have
$\begin{split}\mathcal{L}_{n}(B)-\mathcal{L}_{n}(B_{*})&\leq\lambda_{n}\|\bar{B_{*}}\|_{1}-\lambda_{n}\|\bar{B}_{Pl}\|_{1}\\ &=\lambda_{n}\|\bar{B_{*}}_{\mathcal{S}}\|_{1}-\lambda_{n}\|\bar{B}_{Pl;{\mathcal{S}}}\|_{1}-\lambda_{n}\|\bar{B}_{Pl;{\mathcal{S}^{c}}}\|_{1}\\ &\leq\lambda_{n}\|\tilde{\Delta}_{\mathcal{S}}\|_{1}-\lambda_{n}\|\tilde{\Delta}_{\mathcal{S}^{c}}\|_{1}\end{split}$ (43)
by triangle inequality and $\bar{B_{*}}_{\mathcal{S}^{c}}$, where $\tilde{\Delta}=\bar{B}_{Pl}-\bar{B_{*}}$. We now study the term $\mathcal{L}_{n}(B)-\mathcal{L}_{n}(B_{*})$. Since $\bar{Y}=\bar{X}\bar{B}_{*}+\bar{E}$, we have
$\begin{split}\mathcal{L}_{n}(B)-\mathcal{L}_{n}(B_{*})&=\frac{1}{n}(\bar{Y}-\bar{X}\bar{B}_{Pl})^{T}(\Omega_{n}\otimes I_{n})(\bar{Y}-\bar{X}\bar{B}_{Pl})-\frac{1}{n}(\bar{Y}-\bar{X}\bar{B}_{*})^{T}(\Omega_{n}\otimes I_{n})(\bar{Y}-\bar{X}\bar{B}_{*})\\ &=\frac{1}{n}(\bar{X}\tilde{\Delta})^{T}(\Omega_{n}\otimes I_{n})(\bar{X}\tilde{\Delta})\\ &\indent+\frac{2}{n}(\bar{X}\tilde{\Delta})^{T}(\Omega_{n}\otimes I_{n})\bar{E}.\end{split}$ (44)
Recall that
$v(\eta^{2},\theta)=\frac{1}{\eta^{2}(1+(q-1)\theta)}.$ (45)
By assumption (C3), we get
$\begin{split}\frac{1}{n}(\bar{X}\tilde{\Delta})^{T}(\Omega_{n}\otimes I_{n})(\bar{X}\tilde{\Delta})&\geq\kappa v(\eta^{2}_{n},\theta_{n})\|\tilde{\Delta}\|_{2}^{2}\\ &=\kappa v(\eta^{2}_{*},\theta_{*})\|\tilde{\Delta}\|_{2}^{2}+\kappa v(\eta^{2}_{*},\theta_{*})\left(\frac{v(\eta^{2}_{n},\theta_{n})}{v(\eta^{2}_{*},\theta_{*})}-1\right)\|\tilde{\Delta}\|_{2}^{2}\\ &\geq\kappa v(\eta^{2}_{*},\theta_{*})\|\tilde{\Delta}\|_{2}^{2}-\kappa v(\eta^{2}_{*},\theta_{*})\left|\frac{v(\eta^{2}_{n},\theta_{n})}{v(\eta^{2}_{*},\theta_{*})}-1\right|\|\tilde{\Delta}\|_{2}^{2}\\ &=\kappa v(\eta^{2}_{*},\theta_{*})\left(1-\left|\frac{v(\eta^{2}_{n},\theta_{n})}{v(\eta^{2}_{*},\theta_{*})}-1\right|\right)\|\tilde{\Delta}\|_{2}^{2}\\ &\geq\frac{1}{2}\kappa v(\eta^{2}_{*},\theta_{*})\|\tilde{\Delta}\|_{2}^{2}.\end{split}$ (46)
On the other hand, we have
$\begin{split}\frac{1}{n}(\bar{X}\tilde{\Delta})^{T}(\Omega_{n}\otimes I_{n})\bar{E}&\geq-\frac{1}{n}\|\tilde{\Delta}\|_{1}\left\|\bar{X}^{T}(\Omega_{n}\otimes I_{n})\bar{E}\right\|_{\infty}\\ &=-\frac{1}{n}\|\tilde{\Delta}\|_{1}\left\|\bar{X}^{T}\bar{R}\right\|_{\infty},\end{split}$ (47)
where $(\Omega_{n}\otimes I_{n})\bar{E}=\bar{R}$; $\bar{R}:={\rm vec}(R)$, $R\in\mathbb{R}^{n\times q}$, and $R_{i\cdot}$’s are iid from $N(0,\Omega_{n}\Sigma_{*}\Omega_{n})$, for $i=1,\ldots,n$. Thus, when $\lambda_{n}\geq 4\|\bar{X}^{T}\bar{R}\|_{\infty}/n$, we get
$\begin{split}\frac{2}{n}(\bar{X}\tilde{\Delta})^{T}(\Omega_{n}\otimes I_{n})\bar{E}\geq-\frac{\lambda_{n}}{2}\|\tilde{\Delta}\|_{1}.\end{split}$ (48)
Combining (43)–(48) implies
$\begin{split}\frac{1}{2}\kappa v(\eta^{2}_{*},\theta_{*})\|\tilde{\Delta}\|_{2}^{2}-\frac{\lambda_{n}}{2}\|\tilde{\Delta}\|_{1}\leq\lambda_{n}\|\tilde{\Delta}_{\mathcal{S}}\|_{1}-\lambda_{n}\|\tilde{\Delta}_{\mathcal{S}^{c}}\|_{1},\end{split}$
which in turn yields
$\begin{split}\frac{1}{2}\kappa v(\eta^{2}_{*},\theta_{*})\|\tilde{\Delta}\|_{2}^{2}&\leq\frac{3\lambda_{n}}{2}\|\tilde{\Delta}_{\mathcal{S}}\|_{1}-\frac{\lambda_{n}}{2}\|\tilde{\Delta}_{\mathcal{S}^{c}}\|_{1}\\ &\leq\frac{3\lambda_{n}}{2}\sqrt{s}\|\tilde{\Delta}\|_{2},\end{split}$
and further
$\begin{split}\|\tilde{\Delta}\|_{2}\leq\frac{3\lambda_{n}}{\kappa v_{*}(\eta^{2}_{*},\theta_{*})}\sqrt{s}.\end{split}$ (49)
Note that $R_{\cdot j}$ has iid normal entries with mean zero and variance $r_{n}$, where
$r_{n}$ $\equiv r_{n}(\eta^{2}_{*},\theta_{*},\eta^{2}_{n},\theta_{n})$
$:=\left[(1-\theta_{*})\frac{1+(q-2)\theta_{n}}{1+(q-1)\theta_{n}}-(\theta_{n}-\theta_{*})\frac{1-\theta_{n}}{(1+(q-1)\theta_{n})^{2}}\right]\frac{\eta^{2}_{*}}{(\eta^{2}_{n})^{2}}\frac{1}{(1-\theta_{n})^{2}}.$ (50)
On $\mathcal{C}_{0}$, by (33), $r_{n}$ satisfies
$\begin{split}(\eta^{2}_{*})^{2}(1-\theta_{*})^{2}r_{n}&\leq\left[(1-\theta_{*})+|\theta_{n}-\theta_{*}|\right]\frac{(\eta_{*}^{3})^{3}}{(\eta^{2}_{n})^{2}}\frac{(1-\theta_{*})^{2}}{(1-\theta_{n})^{2}}\\ &\leq 2\eta^{2}_{*}(1-\theta_{*}).\end{split}$
As in the proof of Theorem 2, we define the events
$\mathcal{C}_{j}=\left\{\frac{1}{n}\|X^{T}R_{\cdot j}\|_{\infty}\leq\sqrt{d_{n}\frac{t+2\log p}{n}}\right\},$
$\mathcal{C}=\cap_{j=1}^{q}\mathcal{C}_{j},$
where $d_{n}:=\frac{2}{\eta^{2}_{*}(1-\theta_{*})}$. Since $\mathbb{P}(\mathcal{C}_{j}^{c}|\mathcal{C}_{0})\leq 2\exp(-t/2)$ by Lemma 6.2 of Bühlmann and Van De Geer [2] and Assumption (C1)(C2), and (C4), we have
$\mathbb{P}(\mathcal{C}|\mathcal{C}_{0})=\mathbb{P}(\cap_{j=1}^{q}\mathcal{C}_{j}|\mathcal{C}_{0})\geq 1-\sum_{j=1}^{q}\mathbb{P}(\mathcal{C}_{j}^{c}|\mathcal{C}_{0})\geq 1-2q\exp(-t/2).$
Recall that $\lambda_{n}=8\sqrt{d_{n}\log(pq)/n}$. Thus, again letting $t=4\log q+2\log p$, we have
$\begin{split}\mathbb{P}\left(\lambda_{n}\geq 4\|\bar{X}^{T}\bar{R}\|_{\infty}|\mathcal{C}_{0}\right)&\geq 1-2\exp(-\log(pq)),\end{split}$
which in turn yields,
$\begin{split}\mathbb{P}\left(\lambda_{n}\geq 4\|\bar{X}^{T}\bar{R}\|_{\infty}\right)&\geq(1-2\exp(-\log(pq)))(1-k_{1}\exp(-k_{2}\log(pq)))\\ &\geq 1-(2+k_{1})\exp(-\log(pq)),\end{split}$ (51)
due to Assumption (C4). Combining (51) and (49) completes the proof with $c_{4}=k_{1}+2$, $c_{5}=1$, $c_{6}=24\sqrt{2}$. ∎
Appendix G Simulations for comparison of tuning parameter selection procedure
In this section, we conduct simulations that support the tuning parameter selection procedures introduced in Section 6. Throughout this section, we follow the same data generating procedure used in Section 7 (see (16) and the following paragraph).
G.1 Selection procedures comparison for MRCS, ap.MRCS
We set $s_{1}=s_{2}=0.1$ and $\eta_{i}=1$ for all $i\in\{1,\ldots,q\}$, $(p,q)\in\{(20,50),(50,20)\}$. For MRCS-Or, MRCS (which is excluded in $(p,q)=(50,20)$), and ap.MRCS, we compared the approximate Gaussian validation likelihood minimization [GVLM] to the validation prediction error minimization [VPEM], where the validation prediction error is computed by $\sum_{k=1}^{K}\|Y_{k}-X_{k}\hat{B}_{-k}\|_{F}^{2}$ (see Section 6 for the definition of $K$, $B_{-k}$, $Y_{k}$ and $X_{k}$). For both procedures, we select each tuning parameter with 5 fold cross validation from $\lambda\in\Lambda$, where $\Lambda=\{10^{-4+0.5k}:k=0,1,\ldots,14\}$.
As shown in Figure 1(a) and 1(b), VPEM struggled especially with large $\theta$ values. As opposed to VPEM, GVLM did not suffer the same issue. TPR does not show any difference between VPEM and GVLM but all estimator improves as $\theta$ increases (see Figure 1(d). However, with respect to TNR, GVLM has better TNR results than VPEM (see Figure 1(c)).
(a) $(p,q)=(20,50)$; Model error
(b) $(p,q)=(50,20)$; Model error
(c) $(p,q)=(20,50)$; TNR
(d) $(p,q)=(20,50)$; TPR
Figure 1: Comparison of tuning parameter selection procedures: The setting is from Section 7.1 where $s_{1}=s_{2}=0.1$, $\eta=1,\theta\in\{0,0.5,0.75,0.9,0.95\}$, $(p,q)=(20,50)$ or $(50,20)$, $n=50$. “pre” stands for the prediction loss [VPEM], and “val” denotes the approximate Gaussian validation likelihood used for cross validation [GVLM]. We graphed the model error, ${\rm tr}[(\hat{B}-B)^{T}\Sigma_{X}(\hat{B}-B)]$, TNR and TPR of each procedure with varying $\theta$. The results are based on 50 independent replications.
G.2 Selection procedures comparison for MRGCS, ap.MRGCS
In addition to the tuning parameter selection procedure (15), one might also consider using the estimated parameters from the output of ap.MRGCS, which is given by
$\hat{\lambda}=\mathop{arg\,min}\displaylimits_{\lambda\in\Lambda}\sum_{k=1}^{K}\operatorname{tr}\left\{n^{-1}(Y_{k}-X_{k}\hat{B}_{-k})^{\prime}(Y_{k}-X_{k}\hat{B}_{-k})\hat{\Sigma}_{-k}^{-1}\right\},$
where
$\hat{\Sigma}_{-k}={\rm diag}(\{\hat{\eta}_{-k,i}\}_{i=1}^{q})\left[(1-\hat{\theta}_{-k})I_{q}+\hat{\theta}_{-k}1_{q}1_{q}^{\prime}\right]{\rm diag}(\{\hat{\eta}_{-k,i}\}_{i=1}^{q}).$
In the preceding displays, $\hat{\eta}_{-k,i},\hat{\theta}_{-k}$ are computed from Step 1 of Algorithm 4.1 based on the subsample that excludes the $k$th fold. However, as noted in Section 6, we still suggest using (15) for the optimal tuning parameter selection procedure for MRGCS and ap.MRGCS. We validate this proposal through the following simulations.
We still use the same setting as in Section G.1 except that we used (22) and (23) for $\eta_{i}$’s for each $(p,q)=(50,20),\,(20,50)$. We compared the two approximate Gaussian validation likelihood loss criteria (for each ap.MRCS and ap.MRGCS output) to VPEM for MRCS-Or, MRGCS (it is excluded in $(p,q)=(50,20)$), and ap.MRGCS.
In Figure 2(a) and 2(b), the output of ap.MRGCS used in GVLM was even unstable compared with its VPEM. Generally, the ap.MRCS output outperformed the others for both MRGCS and ap.MRGCS. In TPR (see Figure 2(d)), all procedure improves as $\theta$ increases, but the output of ap.MRCS showed better TPR than that of ap.MRGCS. As opposed to TPR, ap.MRGCS output showed better TNR (see Figure 2(c)). This indicates that ap.MRGCS output tends to provide a sparse solution. However, using ap.MRCS output is preferred after considering all model error, TNR, and TPR.
(a) $(p,q)=(20,50)$; Model error
(b) $(p,q)=(50,20)$; Model error
(c) $(p,q)=(20,50)$; TNR
(d) $(p,q)=(20,50)$; TPR
Figure 2: Comparison of tuning parameter selection procedures: The setting is from Section 7.2 where $s_{1}=s_{2}=0.1$, $\theta\in\{0,0.5,0.75,0.9,0.95\}$, $(p,q)=(20,50)$ or $(50,20)$, $n=50$ (see Section 7.2 for the corresponding $\eta_{i}$ values). “pre” stands for the prediction loss, “val1” denotes the approximate Gaussian validation likelihood procedure using ap.MRCS-based covariance, and “val2” denotes the approximate Gaussian validation likelihood procedure using ap.MRGSC-based covariance. We graphed model error, TNR and TPR of each procedure while varying $\theta$. The results are based on 50 independent replications.
Appendix H Additional simulations
H.1 Supplement to Setting I: Low-dimensional simulation
We follow the same simulation setup as in (16)–(18) (see Section 7.1) except that we use $p=q=20$. The complete results are in Figure 36.
Generally, all our equicorrelation based estimators outperformed lasso estimators as well as MRCE and ap.MRCE. As $\theta$ increases, all our estimators improve. ap.MRGCS slightly suffered $(s_{1},s_{2})=(0.5,0.1)$. This is also a common phenomenon in other simulations regardless of $p,q$ (see Section 7.1, 7.2, and H.2) . Among the competitors, MRCS and ap.MRCS were the best non-oracle methods. MRCE and ap.MRCE performed decently as $s_{1}s_{2}$ increases, especially, when $(s_{1},s_{2})=(0.5,1)$. MRCE and ap.MRCE were the worst two methods in TNR. All joint optimization methods improve in TPR as $\theta$ gets increased.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 3: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 4: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 5: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 6: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,20)$.
H.2 Supplement to Setting II: Equicorrelation of $\Sigma_{*}$ with Heterogeneous and symmetrically distributed $\eta_{i}$’s
We follow the same simulation setup as in Section 7.2 except that we used the following symmetric marginal variances:
$\begin{split}&\eta_{1},\ldots,\eta_{5}=1/\sqrt{3},\,\eta_{6},\ldots,\eta_{10}=1/\sqrt{2},\,\eta_{11},\ldots,\eta_{15}=1/\sqrt[4]{3},\,\eta_{16},\ldots,\eta_{20}=1/\sqrt[4]{2},\\ &\eta_{21},\ldots,\eta_{30}=1,\,\eta_{31},\ldots,\eta_{35}=\sqrt[4]{2},\,\eta_{36},\ldots,\eta_{40}=\sqrt[4]{3},\,\eta_{41},\ldots,\eta_{45}=\sqrt{2},\,\eta_{46},\ldots,\eta_{50}=\sqrt{3},\end{split}$ (52)
when $(p,q)=(20,50)$, and
$\begin{split}&\eta_{1},\ldots,\eta_{4}=1/\sqrt{3},\,\eta_{5},\ldots,\eta_{8}=1/\sqrt{2},\,\eta_{9},\ldots,\eta_{12}=1,\\ &\eta_{13},\ldots,\eta_{16}=\sqrt{2},\,\eta_{17},\ldots,\eta_{20}=\sqrt{3},\end{split}$ (53)
when $(p,q)=(50,20)$. Lastly, when $(p,q)=(80,80)$, we considered
$\begin{split}&\eta_{1},\ldots,\eta_{10}=1/\sqrt{3},\,\eta_{11},\ldots,\eta_{20}=1/\sqrt{2},\,\eta_{21},\ldots,\eta_{30}=1/\sqrt[4]{3},\,\eta_{31},\ldots,\eta_{40}=1/\sqrt[4]{2}\\ &\eta_{41},\ldots,\eta_{50}=\sqrt[4]{2},\,\eta_{51},\ldots,\eta_{60}=\sqrt[4]{3},\,\eta_{61},\ldots,\eta_{70}=\sqrt{2},\,\eta_{71},\ldots,\eta_{80}=\sqrt{3}.\end{split}$ (54)
H.2.1 Results when $(p,q)=(20,50)$
The complete results are in Figure 710. MRGCS was the best non-oracle and ap.MRGCS was the second-best. In addition, MRCS and ap.MRCS performed well, even if they assume compound symmetry. This may be due to the symmetrically distributed variances, $\eta_{i}$ ($i\in\{1,\ldots,q\}$), which are likely to smooth out the heterogeneous variance effect. ap.MRCE performed substantially worse than our equicorrelation based estimators. TNR of ap.MRCE was also significantly worse than the others.
H.2.2 Results when $(p,q)=(50,20)$
The complete results are in Figure 1114. The results are similar to those in Section 7.2.2. ap.MRGCS still struggled with $(s_{1},s_{2})=(0.5,0.1)$, but it was the best non-oracle estimator among the non-oracle procedures. In general, ap.MRGCS and ap.MRCS outperformed the others.
H.2.3 Results when $(p,q)=(80,80)$
The complete results are in Figure 1518. The results are similar to those in Section 7.2.3, except that ap.MRCE showed substantially poor performance compared with our methods. ap.MRGCS again struggled with $(s_{1},s_{2})=(0.5,0.1)$, and it was even outperformed by separate lasso at all $\theta$ values. However, except for this case, ap.MRGCS and ap.MRCS showed the best non-oracle performance.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 7: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 8: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 9: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 10: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 11: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 12: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 13: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 14: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 15: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 16: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 17: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 18: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
H.3 Additional simulations to Section 8.2
We use the same data generating process in Section 8.2 (see (25)) except that we consider different procedure for generating the diagonal matrix $D$ in (25). In this section, we generated diagonals of $D$ through iid draws from a chi-square distribution, which leads to a diagonally dominant $\Sigma_{*}$. The following is the condition numbers for each setting:
  • When $(p,q)=(20,50)$
    • $D$ constructed upon iid $\chi^{2}(1)$ draws: 1336.829 / 420.5613 / 449.145 (setting A2 /B2 / C2 according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid $\chi^{2}(3)$ draws: 71.66033 / 134.6858 / 394.8582 (setting D2 / E2 / F2 according to $w=1/0.5/0.05$)
  • When $(p,q)=(50,20)$
    • $D$ constructed upon iid $\chi^{2}(1)$ draws: 832.5564 / 184.3252 / 180.9868 (setting G2 / H2 / I2 according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid $\chi^{2}(3)$ draws: 45.09728 / 48.67809 / 149.5864 (setting J2 / K2 / L2 according to $w=1/0.5/0.05$)
Again, we note that a $q\times q$ compound symmetry matrix with $q=50$ (resp. $q=20$) and $\theta=0.9$ has condition number 451 (resp. 181).
The complete results are in Table 36, and they are similar to the results in Section 8.2. We use the same sub-setting notation from A2 to L2 as in Section 8.2. The oracle estimator was generally the best, except for G2-3, J2-3, K2-3, where the sparsity is violated and the ridge methods performed the best. However, even when the corruption level is high, our equicorrelation based estimators generally showed comparable performance to lasso methods and frequently were the best non-oracle methods, when $B_{*}$ has sparsity. MRCS and ap.MRCS still showed competitive performance to MRGCS and ap.MRGCS.
Prediction error comparison table: Setting A2–F2
Design OLS LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or MRCS ap.MRCS MRGCS ap.MRGCS
A2-1 105.0715 78.79229 76.45057 84.4663 79.51207 90.18449 72.06984 84.04781 80.67247 81.59314 80.86959
(0.90149) (0.35488) (0.39873) (0.40967) (0.4522) (0.49417) (0.36046) (0.47639) (0.4649) (0.47639) (0.4649)
A2-2 104.62121 87.5176 90.19142 94.172 95.97215 87.86725 80.38446 87.58095 87.33436 87.81213 87.35972
(0.76294) (0.42496) (0.47697) (0.47873) (0.47486) (0.44036) (0.37276) (0.43506) (0.42998) (0.43506) (0.42998)
A2-3 105.02951 99.98116 103.14988 98.95046 100.2181 99.4692 96.4327 102.9058 101.93695 103.28047 102.08946
(0.76602) (0.68208) (0.71519) (0.69429) (0.67904) (0.66999) (0.63352) (0.70861) (0.64167) (0.70861) (0.64167)
B2-1 75.65611 56.90228 55.39336 62.02734 58.17915 63.11375 50.72981 56.4124 55.45088 58.04484 55.66515
(0.90073) (0.35437) (0.40303) (0.39724) (0.42282) (0.47894) (0.29174) (0.45323) (0.29716) (0.45323) (0.29716)
B2-2 73.78413 64.02048 64.55626 68.72112 69.41612 63.25231 55.59376 60.41139 60.19168 59.97346 60.34828
(0.88268) (0.48114) (0.4924) (0.56445) (0.56592) (0.45225) (0.31204) (0.3497) (0.39223) (0.3497) (0.39223)
B2-3 75.05742 71.5248 72.97053 70.28107 71.39136 70.17108 67.49098 70.89491 69.86035 70.5892 69.88786
(0.72402) (0.64157) (0.6741) (0.59533) (0.62579) (0.58586) (0.53642) (0.64858) (0.61208) (0.64858) (0.61208)
C2-1 45.69949 35.36147 33.49663 38.78865 35.49974 33.03231 28.54844 28.90827 28.94331 28.95225 29.13247
(1.11297) (0.41784) (0.39197) (0.50628) (0.47922) (0.34002) (0.27932) (0.28021) (0.28293) (0.28021) (0.28293)
C2-2 47.33275 40.75669 41.27053 44.88513 45.12465 35.05755 30.34744 30.4403 30.43892 30.5317 30.52386
(1.17134) (0.62639) (0.65807) (0.77086) (0.78301) (0.51665) (0.36232) (0.35979) (0.36076) (0.35979) (0.36076)
C2-3 48.03612 43.98291 44.6328 43.81589 44.13011 40.61786 37.09105 38.21206 37.59434 40.20935 39.438
(1.2146) (0.81084) (0.80776) (0.76541) (0.78705) (0.64727) (0.5055) (1.10625) (0.90502) (1.10625) (0.90502)
D2-1 276.13188 198.59586 197.13256 206.97201 200.13757 216.46769 190.67826 199.93753 199.54058 200.67057 200.14572
(1.88367) (0.77651) (0.95993) (0.80589) (0.93312) (1.02253) (0.63568) (0.76923) (0.84938) (0.76923) (0.84938)
D2-2 281.52534 225.09303 229.32895 229.17669 234.8962 225.714 214.27238 220.43282 220.36244 222.32033 221.14247
(2.44137) (1.02092) (1.1545) (0.96211) (1.15356) (1.04621) (0.881) (1.00556) (0.99778) (1.00556) (0.99778)
D2-3 273.4428 250.22042 260.06209 244.35205 249.28644 249.79425 242.98559 263.89819 257.34857 264.38402 257.61919
(1.85887) (0.91029) (0.91327) (0.95593) (1.0345) (0.86324) (0.80546) (1.49318) (1.10038) (1.49318) (1.10038)
E2-1 160.90386 120.88791 116.45597 126.2724 119.84545 133.08762 112.5294 120.74961 119.28996 119.2998 119.62929
(1.12376) (0.65809) (0.70934) (0.75072) (0.76023) (0.79213) (0.55642) (0.59281) (0.62913) (0.59281) (0.62913)
E2-2 161.12426 130.2945 135.11699 138.43316 141.62628 133.91633 123.78783 128.16975 128.14594 128.93577 128.53518
(1.23431) (0.64501) (0.75366) (0.67923) (0.74148) (0.71979) (0.60883) (0.59831) (0.59708) (0.59831) (0.59708)
E2-3 159.54963 151.26322 155.36816 147.18217 150.72091 150.27369 145.08906 154.4475 151.46421 155.31105 152.08549
(1.17964) (0.85091) (0.87853) (0.71254) (0.76721) (0.77189) (0.71307) (1.16565) (0.86897) (1.16565) (0.86897)
F2-1 54.84875 42.29454 40.23873 46.04544 42.35683 41.81673 35.38748 35.80085 35.92453 35.95849 36.22753
(1.11701) (0.5066) (0.46356) (0.53677) (0.50129) (0.37382) (0.32448) (0.32965) (0.33705) (0.32965) (0.33705)
F2-2 54.82139 48.73603 49.24973 53.24709 53.9335 44.51893 38.45547 38.96842 38.82818 39.23644 39.00763
(1.12467) (0.72249) (0.73747) (0.89498) (0.93141) (0.54087) (0.33838) (0.33371) (0.32464) (0.33371) (0.32464)
F2-3 54.87582 53.39429 54.09649 52.83553 53.33211 50.14688 46.50683 48.06138 47.15296 48.65246 47.27893
(1.12495) (0.77594) (0.77092) (0.73876) (0.75908) (0.6493) (0.57486) (0.83316) (0.6785) (0.83316) (0.6785)
Table 3: The prediction error comparison table for each simulation setting in Section H.3 when $(p,q)=(20,50)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the seventh column.
Model error comparison table: Setting A2–F2
Design OLS LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or MRCS ap.MRCS MRGCS ap.MRGCS
A2-1 42.39295 17.26117 14.91861 22.86371 17.90237 28.43493 10.51511 22.33984 19.11891 20.02997 19.31545
(0.70089) (0.27595) (0.32457) (0.33958) (0.37709) (0.39772) (0.27767) (0.51384) (0.34702) (0.4011) (0.39004)
A2-2 42.23756 25.71624 28.42492 32.30901 34.23738 26.05191 18.57594 25.7862 25.53783 26.0215 25.56082
(0.73002) (0.37409) (0.43187) (0.41279) (0.43208) (0.38233) (0.29277) (0.37367) (0.3739) (0.3611) (0.36155)
A2-3 42.65128 37.52372 40.67412 36.61466 37.76556 37.04592 33.97818 40.38988 39.4397 40.77066 39.60936
(0.71445) (0.48534) (0.52458) (0.51837) (0.49354) (0.47818) (0.41658) (0.46421) (0.45246) (0.53596) (0.45182)
B2-1 31.14934 13.29088 11.69225 18.37806 14.47773 19.42273 7.12205 12.78614 11.82619 14.38179 12.05595
(0.68511) (0.2843) (0.30887) (0.31959) (0.34524) (0.4285) (0.22313) (0.34926) (0.27888) (0.35786) (0.25494)
B2-2 29.65723 20.09906 20.58149 24.75082 25.46399 19.28519 11.8882 16.59847 16.41701 16.2451 16.59242
(0.77201) (0.34665) (0.35734) (0.39576) (0.39688) (0.32426) (0.19258) (0.32694) (0.27514) (0.22163) (0.28141)
B2-3 30.39852 27.37935 28.80194 26.13305 27.22649 26.15787 23.36803 26.77497 25.7945 26.45568 25.79595
(0.52269) (0.50579) (0.5272) (0.4656) (0.48522) (0.48074) (0.3862) (0.49644) (0.46687) (0.50936) (0.46745)
C2-1 18.96585 9.03644 7.02547 12.40961 9.00063 6.50119 1.80521 2.17131 2.18127 2.22184 2.38911
(0.92727) (0.39136) (0.36382) (0.47395) (0.42804) (0.18319) (0.03815) (0.03409) (0.06256) (0.03649) (0.06656)
C2-2 19.15577 13.80474 14.26366 17.88023 18.15454 8.10232 3.37173 3.43928 3.43806 3.54561 3.53265
(1.02716) (0.54388) (0.57717) (0.6924) (0.70646) (0.3728) (0.08005) (0.07848) (0.07352) (0.07175) (0.07201)
C2-3 20.26237 17.76817 18.39034 17.55728 17.91135 14.07746 10.28052 11.27033 10.69155 13.05092 12.43701
(1.07382) (0.75564) (0.74578) (0.72232) (0.72971) (0.582) (0.38326) (0.40731) (0.39051) (1.01575) (0.80875)
D2-1 111.27776 36.45598 34.84428 44.65856 37.88761 54.20788 28.62152 37.81118 37.31678 38.51193 37.937
(1.66321) (0.61624) (0.74871) (0.57515) (0.71301) (0.72843) (0.4955) (0.57045) (0.59651) (0.55557) (0.63313)
D2-2 116.09085 61.49551 65.82194 65.64489 71.40972 62.11997 50.92173 56.98089 56.88708 58.8232 57.63959
(2.1362) (0.90658) (1.04345) (0.83411) (0.9927) (0.93692) (0.76843) (0.85082) (0.84472) (0.89514) (0.8681)
D2-3 110.87053 87.1054 96.81336 81.29844 86.20105 86.89881 80.02526 101.0393 94.41737 101.20467 94.77628
(1.73308) (0.90066) (1.00135) (0.9457) (0.99948) (0.8451) (0.792) (1.31973) (1.08012) (1.50165) (1.11487)
E2-1 66.26434 26.82913 22.55788 32.1563 25.91308 38.70073 18.46266 26.49327 25.10216 25.19701 25.40263
(0.84762) (0.4813) (0.55399) (0.54483) (0.57146) (0.60633) (0.31511) (0.57516) (0.46719) (0.36249) (0.4608)
E2-2 66.73458 36.02501 40.82276 43.92905 47.15565 39.59048 29.6172 33.85957 33.84195 34.64423 34.24579
(1.2566) (0.48539) (0.60869) (0.51081) (0.58005) (0.57995) (0.40664) (0.43007) (0.42747) (0.42874) (0.42352)
E2-3 64.83845 57.19454 61.22643 53.00983 56.59714 56.07241 50.89877 59.98199 57.08262 60.85837 57.61517
(1.02699) (0.77782) (0.79812) (0.59635) (0.70367) (0.64918) (0.60097) (0.79877) (0.72041) (1.03405) (0.73279)
F2-1 22.1054 10.10014 8.16875 13.75242 10.18145 9.58014 3.33699 3.75176 3.84514 3.91707 4.14802
(0.999) (0.41701) (0.35764) (0.43543) (0.39115) (0.22971) (0.08664) (0.08877) (0.10269) (0.09189) (0.10806)
F2-2 22.3724 15.31959 15.84472 19.50745 20.20755 11.55726 5.88562 6.37313 6.22846 6.64052 6.38857
(1.08322) (0.48125) (0.5439) (0.62453) (0.66363) (0.41419) (0.1019) (0.13122) (0.12121) (0.13676) (0.11985)
F2-3 22.04484 20.40432 21.06828 19.82113 20.35087 17.13677 13.42993 15.02451 13.99176 15.63436 14.36106
(0.92716) (0.68318) (0.67232) (0.61695) (0.64229) (0.49477) (0.28681) (0.39563) (0.27427) (0.63396) (0.4605)
Table 4: The model error comparison table for each simulation setting in Section H.3 when $(p,q)=(20,50)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the seventh column.
Prediction error comparison table: Setting G2–L2
$\theta$ LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or ap.MRCS ap.MRGCS
G2-1 44.83533 43.46363 53.05025 51.21144 44.5823 33.9537 45.35985 45.90278
(1.1591) (1.40537) (1.24214) (1.59279) (0.95241) (0.89919) (1.18674) (1.20162)
G2-2 64.82553 69.44478 76.63785 81.77781 62.50284 50.17423 64.92464 64.32426
(0.97176) (1.0354) (0.94248) (1.01504) (0.92436) (1.18861) (1.19798) (1.06947)
G2-3 133.45879 153.04183 109.90867 117.93036 139.2495 146.8506 148.6895 141.73388
(2.66461) (2.8973) (1.59821) (1.74162) (2.60436) (8.81637) (4.86074) (2.94612)
H2-1 37.01575 35.93307 43.34182 42.19558 36.34143 26.23449 33.26838 33.92596
(0.85019) (0.93848) (0.91763) (1.08917) (0.8625) (0.55683) (0.78547) (0.67365)
H2-2 50.89406 55.2831 65.83079 69.89328 50.21027 38.05912 46.45042 46.91264
(0.82046) (1.03371) (0.83084) (0.91223) (0.75308) (0.52668) (0.79867) (0.80098)
H2-3 110.38893 124.38963 93.03523 98.32905 118.24035 92.29068 114.19288 109.58705
(1.66942) (2.04462) (1.46699) (1.57249) (2.35693) (2.19205) (2.7504) (1.67005)
I2-1 29.85667 28.80508 37.88529 36.95055 29.57164 16.49538 17.26156 17.94684
(0.65668) (0.71843) (0.88379) (1.03214) (0.6902) (0.23436) (0.26099) (0.28855)
I2-2 40.65305 44.66089 56.80885 60.82274 38.92704 22.12666 24.56793 25.33271
(0.87366) (1.02929) (1.02794) (1.21179) (0.85435) (0.38204) (0.47369) (0.49249)
I2-3 91.69167 104.63603 78.70372 83.83394 91.35578 60.24393 69.3502 71.55799
(1.96659) (2.1685) (1.55065) (1.54102) (1.97122) (1.26672) (1.20634) (1.22631)
J2-1 125.89001 126.61688 131.36489 126.64774 125.87506 114.49484 128.51114 129.81748
(1.8527) (2.26474) (1.8628) (2.13564) (1.81853) (1.95569) (1.52357) (1.61378)
J2-2 170.12318 177.52194 169.78467 180.68469 166.90405 154.63228 165.53654 166.10404
(1.66048) (1.91271) (1.48657) (1.67556) (1.56726) (1.56597) (1.54722) (1.64566)
J2-3 260.05584 299.78893 223.77021 237.8733 261.45428 257.66007 275.7491 273.79175
(2.41941) (3.48501) (1.96252) (2.01791) (2.60148) (3.09796) (2.90973) (2.99102)
K2-1 82.47142 81.63699 88.23276 87.18334 84.45061 72.24252 80.65078 81.07492
(1.08254) (1.4018) (1.24962) (1.45946) (1.42602) (0.88952) (1.24135) (1.2521)
K2-2 107.31975 119.784 121.33131 127.18878 107.78692 99.23878 110.56906 111.31289
(1.01603) (1.3951) (1.04207) (1.24454) (0.9779) (1.03708) (1.02082) (1.05706)
K2-3 191.46974 220.4474 160.92903 169.7814 188.32649 202.65962 187.86521 188.03239
(1.67299) (3.43336) (1.61539) (1.93251) (1.82336) (5.26571) (2.39414) (2.28333)
L2-1 35.42431 34.2385 43.01718 41.86826 35.28308 23.55794 25.15583 25.66002
(0.69618) (0.72038) (0.85946) (0.98601) (0.69019) (0.349) (0.37057) (0.36527)
L2-2 47.33255 51.39827 62.11514 66.57404 45.6146 30.95948 33.07476 33.79567
(0.96358) (1.10066) (0.91844) (1.01125) (0.76398) (0.42458) (0.44855) (0.4878)
L2-3 105.76418 121.86037 87.06077 93.5535 109.15378 81.73934 89.01742 90.88147
(2.28921) (2.68535) (1.50741) (1.70534) (2.59976) (1.34799) (1.89727) (1.89529)
Table 5: The prediction error comparison table for each simulation setting in Section H.3 when $(p,q)=(50,20)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the sixth column.
Model error comparison table: Setting G2–L2
$\theta$ LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or ap.MRCS ap.MRGCS
G2-1 25.14384 23.78569 33.1064 31.34058 24.88036 14.60273 25.71716 26.31923
(1.13407) (1.34304) (1.21714) (1.54674) (0.93052) (0.87037) (1.17192) (1.1905)
G2-2 44.61278 49.3369 56.27542 61.35783 42.29644 30.55962 44.81787 44.19587
(0.89817) (1.00043) (0.88576) (0.9877) (0.84892) (1.12392) (1.12645) (0.98705)
G2-3 111.11682 130.52844 88.65276 96.53391 116.89217 124.53132 126.1663 119.10135
(2.5884) (2.77553) (1.5735) (1.66283) (2.61566) (8.5668) (4.80341) (2.87938)
H2-1 21.95861 20.9488 28.21326 27.16125 21.27333 11.35252 18.2892 18.86563
(0.823) (0.89498) (0.88127) (1.05464) (0.82758) (0.51332) (0.7277) (0.60315)
H2-2 36.05095 40.32027 50.7382 54.92287 35.24728 23.29013 31.44054 31.91841
(0.79906) (0.97442) (0.76463) (0.85724) (0.70676) (0.49518) (0.72526) (0.74804)
H2-3 93.96859 107.78664 77.1775 82.15628 101.65402 76.4224 97.8597 93.25851
(1.50734) (1.8749) (1.25161) (1.3589) (2.16544) (2.11463) (2.62389) (1.44925)
I2-1 19.13619 18.07609 26.87164 26.03099 18.79242 5.87354 6.64095 7.32643
(0.62648) (0.67067) (0.80152) (0.9598) (0.63511) (0.20828) (0.21676) (0.24041)
I2-2 28.87246 32.92022 44.66045 48.68481 27.18182 10.96406 13.25166 14.0125
(0.75068) (0.89493) (0.86671) (1.08192) (0.71677) (0.29269) (0.3523) (0.37053)
I2-3 80.00193 92.9246 67.27094 72.27628 79.58536 48.6636 57.80099 60.1247
(1.85136) (2.02994) (1.44039) (1.42068) (1.82847) (1.04336) (1.03631) (1.09194)
J2-1 54.98627 56.04015 60.76481 56.25554 54.97554 44.10214 57.56517 58.83578
(1.73007) (2.20274) (1.75115) (2.0839) (1.6641) (1.84991) (1.43305) (1.5081)
J2-2 98.1547 105.43109 98.07573 108.58913 94.82636 82.74791 93.74954 94.29603
(1.39359) (1.76736) (1.14373) (1.45523) (1.3107) (1.30263) (1.25026) (1.33995)
J2-3 184.39942 223.20874 149.23673 163.32049 185.62347 182.5674 200.04534 197.98187
(2.10522) (2.87413) (1.55949) (1.7028) (2.28317) (2.94675) (2.71763) (2.72995)
K2-1 41.87521 41.08228 47.69412 46.6846 43.84725 31.72079 40.02228 40.49055
(1.02633) (1.4252) (1.17469) (1.44253) (1.28318) (0.81321) (1.13387) (1.16072)
K2-2 66.0407 78.16363 80.22467 85.91252 66.47505 58.40079 69.28555 69.99709
(0.87697) (1.31751) (0.91312) (1.16308) (0.8628) (0.92193) (0.90648) (0.90859)
K2-3 149.12254 177.60716 119.57849 128.35844 146.01533 160.09242 145.69808 145.89654
(1.44325) (3.14376) (1.33734) (1.63375) (1.55979) (4.91225) (2.21759) (2.11809)
L2-1 21.67108 20.48414 29.25733 28.07226 21.52181 10.12368 11.71929 12.19634
(0.65689) (0.67071) (0.82955) (0.97772) (0.65743) (0.31029) (0.35248) (0.34613)
L2-2 33.68214 37.71081 48.27169 52.5079 32.09719 17.53978 19.71687 20.40054
(0.84221) (0.99064) (0.73503) (0.88085) (0.63973) (0.37087) (0.40564) (0.43919)
L2-3 90.6962 106.13995 72.41254 78.84207 93.63538 66.93323 74.10679 75.99779
(2.18118) (2.37254) (1.32946) (1.53518) (2.31693) (1.25891) (1.7931) (1.7611)
Table 6: The model error comparison table for each simulation setting in Section H.3 when $(p,q)=(50,20)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the sixth column.
Appendix I Tables and figures
I.1 Tables for Section 8.2
Prediction error comparison table: Setting A–F
Design OLS LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or MRCS ap.MRCS MRGCS ap.MRGCS
A-1 421.3023 298.42428 297.6855 306.96871 298.78469 314.64878 265.44132 297.93267 299.79817 299.67561 299.8103
(3.28586) (1.32846) (1.47443) (1.4173) (1.39863) (1.77209) (1.10646) (1.34256) (1.28261) (1.34256) (1.28261)
A-2 421.24545 327.56468 338.57449 333.10994 341.87992 328.97359 295.09134 327.90459 327.65685 330.86734 328.64242
(2.90543) (1.25869) (1.5704) (1.40993) (1.48088) (1.41943) (1.31911) (1.48397) (1.40809) (1.48397) (1.40809)
A-3 423.84581 372.56158 386.54732 360.10005 368.7804 370.65029 347.78963 394.76222 383.01155 393.60105 382.69695
(2.94269) (2.0557) (2.11265) (1.80145) (1.99544) (1.8931) (1.72325) (2.70602) (1.87572) (2.70602) (1.87572)
B-1 238.43803 171.61615 168.99091 178.45972 172.44148 184.72421 151.78365 170.77475 169.45224 171.33406 170.15885
(1.99238) (0.77837) (0.89788) (0.7475) (0.86062) (0.87895) (0.60891) (0.59449) (0.60621) (0.59449) (0.60621)
B-2 231.72801 189.36662 194.28062 197.3021 200.55102 192.46704 167.95685 184.63716 184.72838 185.73553 185.13032
(2.3423) (0.83056) (1.06268) (0.7944) (1.06139) (0.90105) (0.70579) (0.80241) (0.76253) (0.80241) (0.76253)
B-3 234.71038 214.87314 221.11154 207.72906 213.15969 211.66591 202.6108 223.0754 217.47247 222.62461 217.58813
(1.74771) (1.44266) (1.65027) (1.33315) (1.42005) (1.34992) (1.32292) (1.70947) (1.54624) (1.70947) (1.54624)
C-1 61.51457 46.67554 45.16545 51.09038 47.56594 47.68118 39.46289 41.24857 41.41736 41.60935 41.83975
(1.14936) (0.41137) (0.41993) (0.54318) (0.50113) (0.39966) (0.30375) (0.31877) (0.32009) (0.31877) (0.32009)
C-2 63.29166 54.5344 54.65462 58.97623 59.39909 50.45691 43.35193 45.68504 44.94594 46.09739 45.12151
(1.19927) (0.64778) (0.66049) (0.78077) (0.74672) (0.51123) (0.40627) (0.44356) (0.43563) (0.44356) (0.43563)
C-3 64.03305 59.30977 60.07406 58.46108 59.34065 55.4602 52.10512 53.49548 52.51051 53.94183 53.05081
(1.25486) (0.81106) (0.81974) (0.77287) (0.79913) (0.58784) (0.50461) (0.626) (0.57489) (0.626) (0.57489)
D-1 558.74471 390.47356 388.68335 395.06585 389.16316 409.55343 358.98208 385.93684 386.26484 388.81632 387.33796
(3.73806) (1.24105) (1.63307) (1.34186) (1.48051) (1.62612) (1.28938) (1.35932) (1.29872) (1.35932) (1.29872)
D-2 564.13649 424.37231 442.28868 434.68602 442.94447 430.49058 404.62239 429.33388 427.00528 433.22052 428.0325
(5.0482) (1.51057) (1.84653) (1.44811) (1.67828) (1.46774) (1.47361) (1.51177) (1.45592) (1.51177) (1.45592)
D-3 555.82106 483.69323 495.75362 457.10829 472.06041 478.29561 490.9024 508.10263 488.57275 509.09845 486.62817
(3.90354) (1.67152) (1.83794) (1.40863) (1.58048) (1.54378) (2.3392) (3.64916) (1.97457) (3.64916) (1.97457)
E-1 303.01131 216.49803 215.36076 225.8805 217.66458 235.32617 199.25027 216.25975 215.38208 217.45551 216.36837
(1.99873) (0.94551) (1.06092) (1.06224) (1.10995) (1.24057) (0.86844) (0.96294) (0.94547) (0.96294) (0.94547)
E-2 302.62906 243.39502 246.56048 245.26924 251.98864 241.48383 222.12474 235.85209 236.39031 237.44853 237.10805
(2.11844) (1.07497) (1.13222) (0.98032) (1.07024) (1.05926) (0.96153) (1.02713) (1.02652) (1.02713) (1.02652)
E-3 301.59992 271.84761 283.47367 266.58831 270.46679 271.95514 260.94096 287.39082 278.72921 287.5462 279.35487
(1.97572) (1.33779) (1.52899) (1.29148) (1.31735) (1.26651) (1.21011) (2.27853) (1.41849) (2.27853) (1.41849)
F-1 69.1122 52.27008 50.49189 56.974 52.88907 54.57006 45.21246 46.98724 47.15466 47.78788 47.60962
(1.13611) (0.48932) (0.48904) (0.55182) (0.53755) (0.39142) (0.34061) (0.41626) (0.38378) (0.41626) (0.38378)
F-2 69.0662 60.86532 61.14864 65.8587 66.55683 57.72872 49.90487 52.93111 51.62936 52.90665 51.90156
(1.15143) (0.75389) (0.75718) (0.9132) (0.91706) (0.5888) (0.35281) (0.41744) (0.3619) (0.41744) (0.3619)
F-3 68.93688 66.90949 67.81019 65.59747 66.58424 63.20517 60.10471 62.18728 60.80521 61.87549 60.90517
(1.17405) (0.784) (0.79865) (0.75564) (0.77268) (0.6412) (0.60468) (0.70581) (0.68918) (0.70581) (0.68918)
Table 7: The prediction error comparison table for each simulation setting in Section 8.2 when $(p,q)=(20,50)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the seventh column.
Model error comparison table: Setting A–F
Design OLS LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or MRCS ap.MRCS MRGCS ap.MRGCS
A-1 169.49322 49.93963 49.35352 58.52519 50.50883 66.30341 17.82823 49.50422 51.40402 51.2598 51.45167
(2.53288) (0.96146) (1.13484) (1.0516) (1.1362) (1.39676) (0.67809) (0.95861) (1.11572) (0.96324) (1.03735)
A-2 172.55375 78.36706 89.33389 84.15166 92.95695 79.69219 46.0565 78.65624 78.41338 81.62068 79.38073
(2.62797) (0.93968) (1.04166) (1.01617) (1.06284) (0.96694) (0.898) (0.96127) (0.93359) (1.04528) (0.94881)
A-3 173.05229 121.56443 135.33523 109.53369 117.89774 120.13303 97.57773 144.01482 132.24879 142.4011 131.88986
(2.7413) (1.47978) (1.64689) (1.13637) (1.42685) (1.29742) (1.43399) (1.85565) (1.35438) (2.52323) (1.46298)
B-1 99.23479 34.75711 31.93273 41.61723 35.47491 47.55133 14.7796 33.88262 32.61942 34.36268 33.26948
(1.57555) (0.74709) (0.84364) (0.66706) (0.74302) (0.81709) (0.50105) (0.67183) (0.63723) (0.5858) (0.59968)
B-2 94.38538 51.24645 55.9567 59.05804 62.26096 54.24585 30.57363 46.65243 46.73351 47.7272 47.12427
(1.9962) (0.62806) (0.85358) (0.59067) (0.79994) (0.67322) (0.50754) (0.56914) (0.55903) (0.61967) (0.58215)
B-3 96.1279 76.44908 82.68874 69.60995 74.83864 73.39833 64.36683 84.57594 79.18676 84.35722 79.26583
(1.54068) (1.09613) (1.4136) (1.04494) (1.17936) (1.02504) (0.96925) (1.31653) (1.22334) (1.55367) (1.24824)
C-1 25.47427 11.0638 9.43837 15.46791 11.76765 11.84424 3.45578 5.26781 5.44013 5.61142 5.82692
(0.95087) (0.39081) (0.38988) (0.50382) (0.45624) (0.2639) (0.10422) (0.11345) (0.12778) (0.13509) (0.13586)
C-2 25.71332 18.24156 18.29504 22.6413 23.07493 14.17451 7.0494 9.37432 8.62192 9.77303 8.81697
(1.05639) (0.55983) (0.5722) (0.6873) (0.66562) (0.43169) (0.13847) (0.21277) (0.17155) (0.21043) (0.17246)
C-3 26.83281 23.69789 24.44588 22.85877 23.73203 19.61069 15.97679 17.29766 16.36872 17.72032 16.91631
(1.11853) (0.75449) (0.7476) (0.69729) (0.72958) (0.50636) (0.38878) (0.45466) (0.37332) (0.49317) (0.44281)
D-1 224.83755 62.69404 60.66062 67.14954 61.20804 81.3977 31.40809 58.13068 58.45669 61.09928 59.57743
(3.33543) (0.8991) (1.33599) (0.98038) (1.18794) (1.37702) (0.95891) (0.96301) (0.95851) (1.04411) (0.95744)
D-2 233.11746 95.07723 112.66173 104.91901 112.83378 101.05026 75.24154 100.03988 97.73367 103.84379 98.68772
(4.37987) (1.31886) (1.53547) (1.26361) (1.33494) (1.29215) (1.27828) (1.33791) (1.28075) (1.32325) (1.24385)
D-3 227.17133 156.24403 167.9195 130.21511 144.77209 151.14687 164.11705 180.5606 161.64299 181.52586 159.37811
(3.47871) (1.57096) (1.58083) (1.24865) (1.46772) (1.40266) (1.97114) (2.94715) (1.91862) (3.59713) (1.73592)
E-1 124.24667 39.5913 38.56105 49.04118 41.00836 57.99386 22.35643 39.40462 38.52094 40.51292 39.43874
(1.55272) (0.70682) (0.89482) (0.72425) (0.88218) (0.89817) (0.55781) (0.57523) (0.59058) (0.65933) (0.63862)
E-2 124.17102 65.81175 68.87426 67.56181 74.19776 64.07924 45.13431 58.41796 58.95914 59.98033 59.64012
(2.17952) (0.86855) (0.91018) (0.71284) (0.84761) (0.79321) (0.70173) (0.74688) (0.75938) (0.76769) (0.76233)
E-3 122.85455 94.9139 106.74197 90.07125 93.97887 95.09407 83.92015 109.96276 101.41468 110.34212 102.06991
(1.69619) (1.03329) (1.27066) (1.05467) (1.10338) (0.895) (0.91605) (1.31683) (1.0535) (2.0455) (1.10863)
F-1 27.92716 11.73688 10.14497 16.37945 12.42706 13.99748 4.88248 6.60306 6.7363 7.38975 7.14922
(1.03375) (0.38946) (0.37853) (0.45912) (0.43389) (0.30284) (0.15396) (0.18474) (0.18886) (0.24697) (0.18346)
F-2 28.22102 19.13628 19.40333 23.80962 24.51151 16.3342 9.01109 11.98734 10.62209 11.89068 10.85284
(1.10901) (0.51893) (0.56308) (0.63592) (0.64924) (0.44749) (0.15183) (0.23445) (0.19826) (0.23925) (0.20617)
F-3 27.79606 25.51196 26.37553 24.26311 25.26677 21.83446 18.66392 20.67019 19.38073 20.59742 19.55339
(0.97823) (0.68362) (0.67819) (0.63958) (0.65867) (0.4816) (0.31219) (0.43667) (0.32098) (0.41504) (0.34476)
Table 8: The model error comparison table for each simulation setting in Section 8.2 when $(p,q)=(20,50)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the seventh column.
Prediction error comparison table: Setting G–L
$\theta$ LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or ap.MRCS ap.MRGCS
G-1 216.36174 215.52127 216.47546 211.91985 220.51808 187.96523 222.52834 222.66418
(2.86649) (3.70502) (2.76849) (3.3165) (3.39456) (2.78338) (3.5455) (3.41285)
G-2 263.41619 285.02027 264.92923 276.98381 264.15557 240.47493 262.68599 262.68063
(1.98844) (2.71801) (1.6942) (2.10046) (2.09645) (1.82606) (1.98103) (2.0417)
G-3 395.69762 433.34207 345.28576 359.9579 391.58235 396.31235 396.43227 395.78278
(3.07959) (4.7154) (2.39476) (2.55413) (3.56139) (7.71721) (3.28527) (3.34694)
H-1 126.19852 126.8254 132.26305 127.90539 126.01257 110.17097 128.74505 129.11045
(1.93327) (2.26943) (1.9858) (2.15603) (1.61227) (1.78233) (1.58099) (1.52652)
H-2 173.69427 181.23264 174.25862 184.63054 169.0882 151.32558 168.31328 168.1326
(1.86647) (2.30472) (1.37817) (1.81405) (1.73478) (1.41248) (1.65608) (1.63334)
H-3 261.81695 298.17749 225.74933 241.20864 264.54552 267.58151 276.20571 275.65074
(3.00509) (3.82723) (2.24818) (2.53746) (2.63176) (3.43096) (2.95179) (3.07556)
I-1 41.63441 40.1446 49.17791 48.07737 40.77438 30.08706 33.58747 34.25374
(0.81513) (0.75983) (0.91626) (1.08256) (0.78999) (0.49678) (0.53835) (0.54227)
I-2 57.17378 61.94668 72.17713 77.38905 55.84518 41.50233 45.42631 45.84982
(0.88511) (1.08764) (1.0231) (1.13609) (0.85055) (0.64667) (0.77073) (0.75067)
I-3 118.15544 134.63409 98.78048 103.92663 125.73519 97.30096 112.20106 110.56873
(2.06356) (2.70658) (1.40466) (1.48789) (2.45548) (1.71619) (1.71757) (1.82489)
J-1 200.43829 198.97452 199.56624 195.67974 204.83302 163.73197 203.90903 205.80332
(2.57578) (3.29339) (2.15461) (2.81587) (2.65951) (2.43815) (2.46013) (2.44725)
J-2 246.25816 267.34432 251.08902 262.35167 249.31589 213.42151 245.31119 245.24282
(1.89352) (2.70812) (1.77016) (2.42688) (2.13663) (2.8495) (1.8859) (1.87364)
J-3 371.40413 403.65612 318.37477 334.12819 362.58295 353.62652 370.5129 370.72084
(3.26637) (3.70714) (2.58085) (3.20183) (3.10843) (9.63633) (2.88689) (2.76742)
K-1 123.58749 125.00033 128.98656 124.66788 122.2293 96.88944 122.45039 123.40318
(1.76113) (2.37222) (1.58809) (1.66317) (1.65914) (1.47821) (1.64529) (1.76577)
K-2 163.17932 170.86639 164.91992 175.38938 159.10629 134.11352 157.75598 158.11155
(1.50443) (2.16995) (1.32417) (1.88923) (1.45617) (2.08164) (1.40035) (1.43937)
K-3 246.58744 288.1152 214.22073 228.53336 249.75552 227.63059 258.47333 256.43076
(2.32164) (4.69415) (1.86577) (2.7029) (2.63663) (4.96864) (3.22339) (3.05944)
L-1 41.30739 40.14746 48.24352 47.19567 40.32924 27.67641 32.61238 32.8494
(0.78605) (0.91095) (0.91235) (1.03426) (0.74479) (0.42484) (0.53586) (0.4831)
L-2 54.56206 58.65616 68.54698 74.00147 53.09332 37.0229 41.52311 42.38986
(0.91428) (1.06783) (0.8875) (1.09136) (0.76509) (0.4834) (0.56277) (0.58594)
L-3 114.83236 134.3758 95.65601 102.51199 123.16977 88.68501 106.0427 105.54697
(2.08432) (2.43628) (1.43781) (1.68005) (2.70735) (1.4142) (1.87293) (2.03414)
Table 9: The prediction error comparison table for each simulation setting in Section 8.2 when $(p,q)=(50,20)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the sixth column.
Model error comparison table: Setting G–L
$\theta$ LAS LASP c.ridge s.ridge ap.MRCE MRCS-Or ap.MRCS ap.MRGCS
G-1 79.22513 78.44911 79.32835 74.85925 82.873 51.75728 84.70892 85.10121
(2.6801) (3.53157) (2.58437) (3.13839) (3.19515) (2.63026) (3.41663) (3.24324)
G-2 128.43497 149.90534 129.98387 141.3373 129.36042 106.48232 127.91495 127.78541
(1.6269) (2.69274) (1.40837) (1.75774) (1.68617) (1.5894) (1.60924) (1.68485)
G-3 255.07622 289.50646 206.12884 219.84926 250.77437 254.81208 256.07743 255.29228
(2.67378) (4.11213) (2.24085) (2.46772) (3.12734) (7.56453) (2.88895) (2.94745)
H-1 53.30905 53.94632 59.2235 55.06309 53.11528 37.55906 55.81225 56.12208
(1.94869) (2.3202) (1.8695) (2.10927) (1.58724) (1.71661) (1.51831) (1.49101)
H-2 100.77962 107.93918 100.9323 110.91099 96.05866 78.34487 95.12575 94.95246
(1.64842) (1.99683) (1.15675) (1.44178) (1.51918) (1.2776) (1.47348) (1.4557)
H-3 186.61943 221.7338 150.47592 166.05064 189.81258 192.20289 201.54967 201.03727
(2.6937) (3.31687) (1.88186) (2.16387) (2.41599) (3.2429) (2.73932) (2.8289)
I-1 24.90288 23.50231 32.5863 31.48974 24.09418 13.62827 17.1114 17.767
(0.75135) (0.73593) (0.89479) (1.03914) (0.74172) (0.47499) (0.50828) (0.50074)
I-2 39.4262 44.12001 53.81345 59.00468 37.95985 24.13748 27.96343 28.39407
(0.72589) (0.93035) (0.84404) (0.9907) (0.68486) (0.53432) (0.62822) (0.60723)
I-3 100.15207 116.5759 81.16564 86.18903 107.01062 79.39258 93.80389 92.19276
(1.89651) (2.48215) (1.31368) (1.385) (2.40456) (1.64254) (1.49155) (1.60842)
J-1 75.66905 74.35807 75.04052 70.96314 79.72131 39.70002 78.60089 80.43842
(2.34009) (3.06346) (1.9716) (2.64702) (2.48701) (2.29301) (2.25136) (2.22912)
J-2 123.05243 142.89941 126.66689 137.30797 125.71054 89.74062 122.1067 121.70773
(1.45407) (2.26714) (1.34028) (1.84595) (1.67539) (2.62063) (1.47239) (1.43164)
J-3 242.40282 273.05055 190.55096 205.75307 234.32533 226.40655 241.93481 242.32914
(2.46072) (3.10686) (1.8896) (2.32508) (2.55086) (9.36907) (2.29192) (2.16553)
K-1 56.05778 57.35036 61.68682 57.58781 54.53711 29.86232 54.58814 55.61436
(1.6237) (2.21442) (1.44577) (1.61418) (1.49022) (1.43188) (1.52938) (1.64884)
K-2 94.47923 102.03861 97.0119 107.00942 90.88292 66.47393 89.82406 90.11066
(1.23192) (2.0064) (0.99208) (1.61376) (1.1931) (1.93766) (1.18895) (1.17277)
K-3 178.01675 218.52553 146.56115 160.68305 181.12488 159.33957 189.23007 187.32817
(2.00415) (4.17954) (1.50611) (2.29974) (2.38633) (4.90595) (2.93437) (2.69811)
L-1 24.8693 23.563 31.86869 30.74856 23.96931 11.59146 16.40505 16.62412
(0.72842) (0.823) (0.89224) (1.0197) (0.7297) (0.37133) (0.48437) (0.43348)
L-2 38.17885 42.18953 51.93105 57.10291 36.72201 20.95248 25.36067 26.15182
(0.82051) (0.94958) (0.73181) (0.95249) (0.64068) (0.43382) (0.48927) (0.51303)
L-3 97.06229 115.93181 78.21545 84.9697 105.08442 71.35839 88.28659 88.02542
(2.01866) (2.34998) (1.33738) (1.51183) (2.50173) (1.43838) (1.76157) (1.91513)
Table 10: The model error comparison table for each simulation setting in Section 8.2 when $(p,q)=(50,20)$. The values are the mean squared prediction errors averaged over 50 independent replications. The numbers in parentheses are normalized sample standard deviations. The column labels are defined in Section 7 and 8.1. The labels 1, 2, 31231,\,2,\,31 , 2 , 3 indicate the sparsity levels $(s_{1},s_{2})=(0.5,0.5),\,(0.5,1),\,(1,1)$, respectively. Boldface indicates the best non-oracle model. The oracle estimator is on the sixth column.
I.2 Additional plots for Section 7.1
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 19: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 20: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 21: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 22: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 23: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 24: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 25: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 26: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 27: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 28: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 29: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 30: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
I.3 Additional plots for Section 7.2
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 31: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 32: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 33: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 34: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(20,50)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 35: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 36: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 37: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 38: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(50,20)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 39: Prediction plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 40: Model error plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 41: TNR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
(a) $(0.1,0.1)$
(b) $(0.5,0.1)$
(c) $(0.1,0.5)$
(d) $(0.5,0.5)$
(e) $(0.1,1)$
(f) $(0.5,1)$
Figure 42: TPR plots: Two numbers in each bracket stand for $(s_{1},s_{2})$ when $(p,q)=(80,80)$.
I.4 Additional plots for Section 8.1
(a) Setting A1
(b) Setting A2
(c) Setting A3
(d) Setting B1
(e) Setting B2
(f) Setting B3
(g) Setting C1
(h) Setting C2
(i) Setting C3
Figure 43: Prediction plots for each setting in Section 8.1.
(a) Setting A1
(b) Setting A2
(c) Setting A3
(d) Setting B1
(e) Setting B2
(f) Setting B3
(g) Setting C1
(h) Setting c2
(i) Setting A3
Figure 44: Model error plots for each setting in Section 8.1.